### v3

# Proper scoring rules for estimation and forecast evaluation

Abstract
Proper scoring rules have been a subject of growing interest in recent years, not only as tools for evaluation of probabilistic forecasts but also as methods for estimating probability distributions. In this article, we review the mathematical foundations of proper scoring rules including general characterization results and important families of scoring rules. We discuss their role in statistics and machine learning for estimation and forecast evaluation. Furthermore, we comment on interesting developments of their usage in applications.
1 Introduction
In recent years, proper scoring rules have emerged as a powerful general approach for estimating probability distributions. In addition to significantly expanding the range of modeling techniques that can be applied in practice, this has also substantially broadened the conceptual understanding of estimation methods. Originally, proper scoring rules were conceived in meteorology as summary statistics for describing the performance of probabilistic forecasts (Murphy and Winkler, 1984), but they also play an important role in economics as tools for belief elicitation (Schotter and Trevino, 2014).
A probabilistic forecast is a probability distribution over the space of the possible outcomes of the future event that is stated by the forecaster. The simplest and most popular case of probabilistic forecasts arises when the outcome is binary, so the probabilistic forecast reduces to issuing a predictive probability of success. Brier (1950) was the first to consider the problem of devising a scoring rule which could not be “played” by a dishonest forecasting agent. He introduced the quadratic scoring rule and showed that it incentivizes a forecasting agent to state his most accurate probability estimate when faced with uncertainty. McCarthy (1956)111better known as the designer of the Lisp programming language. was the first to clearly state the concept of propriety as it is understood today and stated as Equation 1.
Consider a forecasting tournament which consists of a forecasting agent issuing a probabilistic forecast $\mathbb{P}$ for a random variable $Y$ and then receiving a penalty $S(\mathbb{P},y)$ according to the forecast $\mathbb{P}$ and the realization $y$ of $Y$. If the forecasting agent is rational and believes that $Y\sim\mathbb{Q}$ for some $\mathbb{Q}$, then it would issue the forecast $\mathbb{P}$ so as to mimimize the expected value $S(\mathbb{P},\mathbb{Q})=\mathbb{E}_{Y\sim\mathbb{Q}}[S(\mathbb{P},Y)]$ of its penalty. In order to compel the agent to reveal its true belief $\mathbb{Q}$, we must choose $S$ such that $S(\mathbb{P},\mathbb{Q})$ is minimized at $\mathbb{P}=\mathbb{Q}$. It follows that we must have
$S(\mathbb{Q},\mathbb{Q})\leq S(\mathbb{P},\mathbb{Q})$ (1)
for every $\mathbb{P},\mathbb{Q}$ (with equality if and only if $\mathbb{P}=\mathbb{Q}$). In other words, $S$ should be a (strictly) proper scoring rule, see Definition 4.
Example 1 (Brier score).
For a binary outcome $y\in\{0,1\}$, the Brier or quadratic score
$S_{\mathrm{Brier}}(\mathbb{P},y)=(p-y)^{2},$
where $p$ is the predicted probability for the event $\{y=1\}$, is a strictly proper scoring rule (Brier, 1950).
Example 2 (Logarithmic score).
The logarithmic score is given by
$S_{\mathrm{log}}(\mathbb{P},y)=-\log p(y),$
where $p$ is the density of $\mathbb{P}$. It is also known as the ignorance score in meteorology and is a strictly proper scoring rule (Good, 1952), see Example 5 for details.
Example 3 (Continuous ranked probability score (CRPS)).
For a real-valued outcome $y\in\mathbb{R}$, the continuous ranked probability score (CRPS) is defined as
$\operatorname{CRPS}(\mathbb{P},y)=\int\left(F_{\mathbb{P}}(x)-\mathbf{1}\{y\leq x\}\right)^{2}\,\mathrm{d}x,$
where $F_{\mathbb{P}}$ is the cumulative distribution function (CDF) of $\mathbb{P}$ (Matheson and Winkler, 1976). It is a strictly proper scoring rule. See Example 7 for details.
During the 1970s, proper scoring rules were developed into a principled approach for the assessment of probabilistic forecasts with applications in meteorology, economics and psychology (Murphy and Winkler, 1970) and as a foundation for the theory of subjective probability (Savage, 1971). McCarthy (1956) produced a general characterization of proper scoring rules for finite outcome spaces. His characterization was later extended by Hendrickson and Buehler (1971) to absolutely continuous measures on the real line, which was followed by a complete characterization by Gneiting and Raftery (2007) for general measurable spaces, see Theorem 12.
Proper scoring rules have long been recognized as elicitation and aggregation mechanisms for subjective opinion in economics and other social sciences (Jensen and Peterson, 1973). For example, the spherical scoring originates in psychology (Roby, 1964). Tetlock (2005) provides a popular account of studies on the accuracy of probabilistic forecasts concerning political events by subject-matter experts. Another interesting application is the crowd-sourcing of forecasts from multiple human subjects via forecasting tournaments such as the Good Judgement Open and Metaculus, which evaluate forecasts using the Brier and logarithmic scores, respectively (GJOpen, 2025; Metaculus, 2025). A fascinating analysis of the literature on proper scoring rules in applications from 1950 to 2015 is given by Carvalho (2016).
The extant literature contains both of the possible conventions for proper scoring rules: one which requires $S(\mathbb{P},\mathbb{Q})$ to be minimized at $\mathbb{P}=\mathbb{Q}$ and the other requiring it to be maximized instead. The influential paper by Gneiting and Raftery (2007) assumes that better forecasts correspond to higher scoring rule values, but most of the more recent literature has adopted the convention in Equation 1. With a negative orientation, scoring rules should be called loss functions, but it is not in our power to change the naming convention.
Some parts of the literature order the arguments of proper scoring rules differently, that is, it would be $S(y,\mathbb{P})$ instead of $S(\mathbb{P},y)$. We mention this since it may cause confusion when dealing with expected scores $S(\mathbb{P},\mathbb{Q})$. “Scoring rule” is sometimes abbreviated to “score”, which may cause confusion, since “score” typically refers to the gradient of the logarithmic density in statistics, machine learning and econometrics. However, for some classes of proper scoring rules such as kernel scores (Section 3.1) the term has now become established.
1.1 Proper scoring rules in estimation
Proper scoring rules present a general and nuanced approach to the problem of estimating probability distributions, which allows for taking into account computational and statistical considerations relevant to the application.
If we think of $\mathbb{Q}$ as the data distribution and $\mathbb{P}$ as a model distribution, then Equation 1 can be interpreted as saying that the expectation $S(\cdot,\mathbb{Q})=\mathbb{E}_{Y\sim\mathbb{Q}}[S(\cdot,Y)]$ is minimized precisely when the model and data distributions are equal. It is reasonable to expect that given an iid sample $\smash{\{Y_{j}\}_{j=1}^{n}}$ from $\mathbb{Q}$ with empirical distribution ^$\hat{\mathbb{Q}}_{n}$, the minimizer of the empirical mean $S(\cdot,\hat{\mathbb{Q}}_{n})=(1/n)\sum_{j=1}^{n}S(\cdot,Y_{j})$ would give reasonably good estimates of $\mathbb{Q}$ under appropriate regularity conditions. Scoring rule minimization can thus be employed for estimating probability distributions.
For example, let $\{\mathbb{P}_{\theta}\}_{\theta\in\Theta}$ be a parametric family of distributions and assume that $\mathbb{Q}=\mathbb{P}_{\theta_{\ast}}$ for some $\theta_{\ast}\in\Theta$. Consider the estimator
$\hat{\theta}=\operatorname*{arg\,min}_{\theta\in\Theta}\left[\frac{1}{n}\sum_{j=1}^{n}S(\mathbb{P}_{\theta},Y_{j})\right]=\operatorname*{arg\,min}_{\theta\in\Theta}S(\mathbb{P}_{\theta},\hat{\mathbb{Q}}_{n}),$
where $S$ is a strictly proper scoring rule. This is known as minimum scoring rule inference, of which the classical maximum likelihood inference is a special case where $S=S_{\mathrm{log}}$ is the logarithmic score of Example 2 (Dawid et al., 2016). Strict propriety ensures that if $\hat{\theta}$ was minimizing the limit of $S(\mathbb{P}_{\theta},\hat{\mathbb{Q}}_{n})$ instead, which is the expectation $S(\mathbb{P}_{\theta},\mathbb{P}_{\theta^{\ast}})=S(\mathbb{P}_{\theta},\mathbb{Q})$, then $\hat{\theta}=\theta_{\ast}$. Propriety can thus be thought of as the population counterpart of consistency for M-estimators of probability distributions (Huber and Ronchetti, 2009).
Minimization of strictly proper scoring rules can also be used to estimate conditional distributions. Let $\{(X_{j},Y_{j})\}_{j=1}^{n}$ be an iid sample of covariates and responses with $(X_{j},Y_{j})\in\mathcal{X}\times\mathcal{Y}$, and let $\{\mathbb{P}_{\theta,x}:(\theta,x)\in\Theta\times\mathcal{X}\}$ be a parametric family of distributions on $\mathcal{Y}$ that model the conditional distributions $\mathbb{P}_{Y|X=x}$ of the response $Y$ given covariate $X=x$. We consider
$\hat{\theta}=\operatorname*{arg\,min}_{\theta\in\Theta}\left[\frac{1}{n}\sum_{j=1}^{n}S(\mathbb{P}_{\theta,X_{j}},Y_{j})\right]$ (2)
which yields the estimator $\smash{\mathbb{P}_{\hat{\theta},x}}$ of $\mathbb{P}_{Y|X=x}$ for every $x\in\mathcal{X}$. This approach estimates the conditional distribution simultaneously for every quantile and covariate, unlike some classical methods such as quantile regression. In this sense, strictly proper scoring rules play the same role for estimating conditional distributions $\mathbb{P}_{Y|X=x}$, that the squared error loss plays for the conditional expectation $\mathbb{E}[Y|X=x]$. Moreover, in a certain sense, these are the only loss functions which can consistently estimate conditional distributions (see Theorem 8).
Although every proper scoring rule can identify the true model or the conditional distribution as its minimum in the population sense, they exhibit different inductive biases or regularization effects in finite samples. Furthermore, the computational tractability varies substantially depending on the chosen scoring rule, see Section 5.
1.2 Proper scoring rules in forecast evaluation
Gneiting and Katzfuss (2014) have argued that forecasts should be probabilistic, since future outcomes are uncertain and this uncertainty ought to be quantified for better decision making. Probabilistic forecasts are routinely issued in more and more application domains including but not limited to weather and climate prediction, flood risk prediction, renewable energy forecasts, financial risk management, epidemiological prediction, and preventive medicine. Given competing forecasters or forecasting methods, the predictive power is compared with a proper scoring rule $S$. Suppose that there is data $(\mathbb{P}_{i},\mathbb{P}^{\prime}_{i},y_{i})_{i=1}^{n}$ of competing probabilistic forecasts $\mathbb{P}_{i}$, $\mathbb{P}_{i}^{\prime}$ and corresponding verifying observations $y_{i}$ for each forecasting instance $i$ (for example, each day for daily predictions). Then, one compares the average realized scores
$\frac{1}{n}\sum_{i=1}^{n}S(\mathbb{P}_{i},y_{i})\quad\text{and}\quad\frac{1}{n}\sum_{i=1}^{n}S(\mathbb{P}_{i}^{\prime},y_{i}),$ (3)
and gives preference to the forecaster with the lower score. Justification of this approach, even in the absence of stationarity of forecast-observation pairs, can be found in Modeste et al. (2023). Inference methods for differences in predictive performance are discussed by Diebold and Mariano (2002); Giacomini and White (2006); Lai et al. (2011); Henzi and Ziegel (2022); Choe and Ramdas (2024) but this list is not comprehensive. The mentioned works focus on assessing predictive performance comparatively between two forecasters based on proper scoring rules. A comprehensive review of inference methods for assessing predictive performance in relative and absolute terms goes beyond the scope of this paper.
Clearly, the empirical scores in Equation 3 can be also be used as an evaluation metric (validation or testing loss) for comparing the performance of different estimates.
While Equation 1 ensures that the optimal forecaster (or true model) will dominate all other forecasters (models or parameters) with regards to the comparison in Equation 3 in the long term, different proper scoring rules will rank competing imperfect forecasters differently even for infinite amounts of data, see for example Patton (2020). While this may seem as a disadvantage of ranking forecasts with respect to proper scoring rules, it is not surprising and there there is no alternative. Therefore, the scoring rule must be chosen carefully for the comparison of empirical scores to be relevant to the application at hand, since it determines when one imperfect probabilistic forecast should be preferred over another. As a result, there are numerous proposals to tailor proper scoring rules such that they are discriminative with respect to certain regions of interest of the predictive distributions and outcomes, see Section 5.
1.3 Structure of the paper
In Section 2, we formally introduce proper scoring rules, discuss characterizations and interpret their geometric properties. Section 3 considers the most important families of proper scoring rules such as kernel scores, local scoring rules and $f$-scores. Scoring rule decompositions are briefly reviewed in Section 4. The use of proper scoring rules in estimation and forecast evaluation is discussed in Section 5, where we consider aspects of computational tractability and the choice of scoring rules for evaluation. The paper concludes with Section 6, where we mention some limitations and extensions of proper scoring rules. All proofs of formal statements in the paper can be found in the appendix.
2 Proper Scoring Rules
Let $\mathcal{Y}$ be a set, $\mathcal{F}$ be a $\sigma$-algebra on $\mathcal{Y}$ and $\mathcal{P}$ denote a convex set of probability measures on the measure space $(\mathcal{Y},\mathcal{F})$. Let $\overline{\mathbb{R}}=\mathbb{R}\cup\{-\infty,+\infty\}$ be the extended real line. We refer to $\mathcal{Y}$ as outcome space. When $\mathcal{Y}=\mathbb{R}^{d}$ for some $d\geq 1$, we assume that $\mathcal{F}$ is the Borel $\sigma$-algebra.
2.1 Definition
A scoring rule is a function
$S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}},\quad(\mathbb{P},y)\mapsto S(\mathbb{P},y)$
such that $y\mapsto S(\mathbb{P},y)$ is $\mathcal{P}$-quasi-integrable for every $\mathbb{P}\in\mathcal{P}$, that is, the integral $S(\mathbb{P},\mathbb{Q})=\int S(\mathbb{P},y)\,\mathrm{d}\mathbb{Q}(y)$ exists for all $\mathbb{P},\mathbb{Q}\in\mathcal{P}$ but is not necessarily finite. Since we are interested in minimizing $S(\mathbb{P},\mathbb{Q})$ over $\mathbb{P}$, multiplication by a positive constant or addition of a function of $y$ is inconsequential. Thus, two scoring rules $S_{1}$ and $S_{2}$ are said to be equivalent if for all $y\in\mathcal{Y}$ and $\mathbb{P}\in\mathcal{P}$, $S_{1}(\mathbb{P},y)=\alpha S_{2}(\mathbb{P},y)+l(y)$ for some $\alpha>0$ and $\mathcal{P}$-integrable $l:\mathcal{Y}\to\overline{\mathbb{R}}$, and strongly equivalent if, in addition, $\alpha=1$.
Definition 4 (Proper Scoring Rules).
A scoring rule $S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ is called proper if
$S(\mathbb{Q},\mathbb{Q})\leq S(\mathbb{P},\mathbb{Q})$ (4)
for every $\mathbb{P},\mathbb{Q}\in\mathcal{P}$ and strictly proper if Equation 4 holds with equality if and only if $\mathbb{P}=\mathbb{Q}$.
We have defined (strict) propriety of a scoring rule with regard to all measures in $\mathcal{P}$. Since there could be cases when $S$ is a scoring rule naturally defined on $\mathcal{P}\times\mathcal{Y}$ but it is only proper when restricted to some subdomain $\mathcal{P}^{\prime}\times\mathcal{Y}\subseteq\mathcal{P}\times\mathcal{Y}$, it is important that the domain of a proper scoring rule is clearly defined. Gneiting and Raftery (2007) account for this by defining propriety relative to some $\mathcal{P}^{\prime}\subseteq\mathcal{P}$. For reasons of conciseness, we prefer the formulation in Definition 4.
Example 5 (Logarithmic score).
Let $\mu$ be a $\sigma$-finite measure on $\mathcal{Y}$ and $\mathcal{P}$ the set of measures which are absolutely continuous with respect to $\mu$. For example, if $\mathcal{Y}=\mathbb{R}^{d}$ and $\mu$ is the Lebesgue measure, we consider absolutely continuous distributions on $\mathbb{R}^{d}$; if $\mathcal{Y}$ is finite and $\mu$ is the counting measure, we have categorical random variables. The logarithmic score $S_{\mathrm{log}}:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ with $S_{\mathrm{log}}(\mathbb{P},y)=-\log p(y)$, where $p$ is the density of $\mathbb{P}$ with respect to $\mu$, is a strictly proper scoring rule (Good, 1952). The logarithmic score is an example of a local scoring rule, see Section 3.2 for further details. Minimizing the logarithmic score is equivalent to the well-known maximum likelihood principle, which is the most popular method for estimating probability distributions and arguably, the foundation of classical statistical inference.
Example 6 (Hyvärinen score).
[…]
$S_{\mathrm{Hyv\ddot{a}rinen}}(\mathbb{P},\mathbf{y})=\Delta_{\mathbf{y}}\log p(\mathbf{y})+\frac{1}{2}\|\nabla_{\mathbf{y}}\log p(\mathbf{y})\|^{2},$
where $\nabla_{\mathbf{y}}f$ and $\Delta_{\mathbf{y}}f$ denote the gradient $(\partial f/\partial y_{1},\dots,\partial f/\partial y_{d})$ and Laplacian $\sum_{j=1}^{d}\partial^{2}f/\partial y_{j}^{2}$ of $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ with respect to $\mathbf{y}=(y_{1},\dots,y_{d})$, and $p$ is the density of $\mathbb{P}$, is a strictly proper scoring rule (Parry et al., 2012). Minimizing the Hyvärinen score is equivalent to score matching, which is the basis of the score-based modeling techniques behind many of the recent advances in machine learning.
Example 7 (Continuous ranked probability score (CRPS)).
Let $\mathcal{Y}=\mathbb{R}$ and $\mathcal{P}$ be the set of distributions with finite mean. The continuous ranked probability score $\operatorname{CRPS}:\mathcal{P}\to\overline{\mathbb{R}}$ given by $\operatorname{CRPS}(\mathbb{P},y)=\int\left(F_{\mathbb{P}}(x)-\mathbf{1}\{y\leq x\}\right)^{2}\,\mathrm{d}x$, where $F_{\mathbb{P}}$ is the CDF of $\mathbb{P}$, is a strictly proper scoring rule (Matheson and Winkler, 1976). It has a compelling interpretation as the $L^{2}$ distance between the CDFs of $\mathbb{P}$ and the forecast $y$. Alternatively, it can also be seen as the Brier score for the binary event $\mathbf{1}{\{Y\leq x\}}$ integrated over $\mathbb{R}$. Furthermore, it has the equivalent representation
$\operatorname{CRPS}(\mathbb{P},y)=\int|x-y|\,\mathrm{d}\mathbb{P}(x)-\frac{1}{2}\iint|x-x^{\prime}|\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{P}(x^{\prime}),$ (5)
see for example Baringhaus and Franz (2004, Lemma 2.1). The representation in Equation 5 identifies the CRPS as a kernel score, see Section 3.1. It is of great significance in meteorology, where it is the most popular scoring rule for real-valued variables.
While these examples consider rather standard outcome domains $\mathcal{Y}$, proper scoring rules have also been extended and applied to exotic domains such as distributions on the circle (Gneiting and Raftery, 2007), sphere (Takasu et al., 2018), general Riemannian manifolds (Mardia et al., 2016), functional data or stochastic processes (Hayati et al., 2024), point processes (Brehmer et al., 2024) and discrete or continuous trajectories (Bonnier and Oberhauser, 2024).
With the goal of forecast evaluation in mind, the notion of propriety at Equation 4 is easily motivated, compare Equation 3. However, the notion of propriety also appears naturally from the perspective of (distributional) regression, compare Equation 2. The following result has previously appeared in Tsyplakov (2011); Holzmann and Eulert (2014). It is analogous to a result of Banerjee et al. (2005) on conditional expectation and Bregman loss functions.
Theorem 8.
Let $\mathcal{Y}$ be a Polish space and $\mathcal{X}$ be some measurable covariate space. A scoring rule $S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ is (strictly) proper if and only if for every pair of random variables $(X,Y)\in\mathcal{X}\times\mathcal{Y}$ such that the conditional distributions $\mathbb{P}_{Y|X=x}$ are in $\mathcal{P}$, and every family $\{\mathbb{P}^{x}:x\in\mathcal{X}\}\subseteq\mathcal{P}$, the expectation $\mathbb{E}\left[S(\mathbb{P}^{X},Y)\right]$ is minimized if (and only if) $\mathbb{P}^{x}=\mathbb{P}_{Y|X=x}$ for $x\in\mathcal{X}$ $\mathbb{P}_{X}$-almost surely, where $\mathbb{P}_{X}$ is the marginal distribution of $X$.
Mathematically, the concept of proper scoring rules is an extension of the concept of Bregman loss functions to probability distributions (Ovcharov, 2018). These are distance-like functions on Euclidean spaces used in the optimization literature to adapt general methods such as proximal gradient descent to the geometry of specific problems (Teboulle, 2018). Importantly, many boosting algorithms such as AdaBoost are known to be instances of one such method known as mirror descent (Collins et al., 2002; Beja-Battais, 2023).
To every scoring rule $S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$, we associate an entropy $H:\mathcal{P}\to\overline{\mathbb{R}}$ and a divergence $d:\mathcal{P}\times\mathcal{P}\to\overline{\mathbb{R}}$ given by
$H(\mathbb{P})=S(\mathbb{P},\mathbb{P})=\int S(\mathbb{P},y)\,\mathrm{d}\mathbb{P}(y)\qquad\mbox{ and }\qquad d(\mathbb{P},\mathbb{Q})=S(\mathbb{P},\mathbb{Q})-H(\mathbb{Q}).$ (6)
For a (strictly) proper scoring rule, the entropy $H$ is a (strictly) concave function on $\mathcal{P}$, since, by propriety for all $\lambda\in(0,1)$,
$H(\lambda\mathbb{P}+(1-\lambda)\mathbb{Q})=\lambda S(\lambda\mathbb{P}+(1-\lambda)\mathbb{Q},\mathbb{P})+(1-\lambda)S(\lambda\mathbb{P}+(1-\lambda)\mathbb{Q},\mathbb{Q})\\ \geq\lambda S(\mathbb{P},\mathbb{P})+(1-\lambda)S(\mathbb{Q},\mathbb{Q})=\lambda H(\mathbb{P})+(1-\lambda)H(\mathbb{Q}).$
The divergence satisfies $d(\mathbb{P},\mathbb{Q})\geq 0$ for $\mathbb{P},\mathbb{Q}\in\mathcal{P}$ with equality if (and only if) $\mathbb{P}=\mathbb{Q}$. The divergence is generally not a metric on $\mathcal{P}$ and, in fact, it does not even need to be symmetric in its arguments. However, for the important class of kernel scores, which includes the CRPS, symmetry holds and the square root of the divergence is a metric, see Example 11 and Section 3.1.
Example 9 (Logarithmic score).
The logarithmic score has entropy
$H(\mathbb{P})=-\int p(x)\log p(x)\,\mathrm{d}\mu(x),$
which is the Shannon entropy. The associated divergence $d(\mathbb{P},\mathbb{Q})=\int q(y)\log(q(y)/p(y))\,\mathrm{d}\mu(y)=D_{\mathrm{KL}}(\mathbb{Q}||\mathbb{P})$ is the Kullback-Leibler divergence.
Example 10 (Hyvärinen score).
The Hyvärinen score has entropy
$H(\mathbb{P})=-\tfrac{1}{2}\int\|\nabla\log p(\mathbf{x})\|^{2}\,p(\mathbf{x})\,\mathrm{d}\mathbf{x},$
see Remark 1 in the appendix for a derivation. The associated divergence $d(\mathbb{P},\mathbb{Q})=\tfrac{1}{2}\int\|\nabla\log p(\mathbf{y})-\nabla\log q(\mathbf{y})\|^{2}\,q(\mathbf{y})\,\mathrm{d}\mathbf{y}$ is sometimes called score matching distance.
Example 11 (CRPS).
The CRPS has entropy
$H(\mathbb{P})=\int F_{\mathbb{P}}(x)(1-F_{\mathbb{P}}(x))\,\mathrm{d}x$
and a symmetric divergence $d(\mathbb{P},\mathbb{Q})=\int(F_{\mathbb{P}}(y)-F_{\mathbb{Q}}(y))^{2}\,\mathrm{d}y$, whose square root is a metric, compare Example 17.
An alternative popular approach to comparing probability distributions is provided by the family of $\phi$-divergences (also known as $f$-divergences) such as the Hellinger and total variation distances. It should be noted that with the sole exception of the Kullback-Leibler divergence, these do not correspond to proper scoring rules (Csiszár, 1991), although they have been shown to be intimately related (Mohamed and Lakshminarayanan, 2017; Gao et al., 2020).
Proper scoring rules have been studied from functional analysis (Ovcharov, 2015), convex analysis (Williamson and Cranko, 2023) and information geometry (Dawid, 2007) perspectives.
2.2 Characterization
Proper scoring rules can be elegantly characterized in terms of their entropy under minimal assumptions. A scoring rule $S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ is called regular if $H(\mathbb{P})=S(\mathbb{P},\mathbb{P})$ is finite and $S(\mathbb{P},\mathbb{Q})>-\infty$ for every $\mathbb{P},\mathbb{Q}\in\mathcal{P}$ with $\mathbb{P}\neq\mathbb{Q}$.
Theorem 12 (Gneiting and Raftery, 2007).
A regular scoring rule $S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ is (strictly) proper if and only if there is a (strictly) concave function $H:\mathcal{P}\to\mathbb{R}$ such that
$S(\mathbb{P},y)=H(\mathbb{P})+\langle h_{\mathbb{P}},\delta_{y}-\mathbb{P}\rangle=H(\mathbb{P})+h_{\mathbb{P}}(y)-\int h_{\mathbb{P}}(y)\,\mathrm{d}\mathbb{P}(y)$ (7)
for every $\mathbb{P}\in\mathcal{P}$ and $y\in\mathcal{Y}$, where $h_{\mathbb{P}}$ is a supergradient of $H$ at $\mathbb{P}$.
In Equation 7, we use the shorthand $\langle f,\mathbb{P}\rangle$ for the integral $\int f\,\mathrm{d}\mathbb{P}$ of a $\mathbb{P}$-integrable function $f$ on $\mathcal{Y}$, and $\delta_{y}$ denotes the Dirac measure at the point $y$. Furthermore, a supergradient of a concave function $H\colon\mathcal{P}\to\overline{\mathbb{R}}$ at $\mathbb{P}$ is a function $h_{\mathbb{P}}\colon\mathcal{Y}\to\overline{\mathbb{R}}$ that is $\mathbb{Q}$-integrable and
$H(\mathbb{P})+\langle h_{\mathbb{P}},\mathbb{Q}-\mathbb{P}\rangle\geq H(\mathbb{Q})$ (8)
for all $\mathbb{Q}\in\mathcal{P}$. The function $H$ in Equation 7 indeed coincides with the entropy of $S$ defined in Equation 6. A regular proper scoring rule is thus determined by its entropy $H$ and the family $\{h_{\mathbb{P}}:\mathbb{P}\in\mathcal{P}\}$ of chosen supergradients. Theorem 12 is general but also somewhat abstract since supergradients of concave functions on convex sets of probability measures may not be directly intuitive. However, in the following, we will showcase the strength and direct applicability of this result in several instances.
Theorem 12 simplifies if the outcome space $\mathcal{Y}=\{y_{1},\dots,y_{n}\}$ is finite with $n$ elements. Then, the probability measures $\mathbb{P}$ on $\mathcal{Y}$ can be identified with vectors $\mathbf{p}=(p_{j})_{j=1}^{n}$ in the unit simplex $\mathcal{S}_{n}=\{(p_{j})_{j=1}^{n}\in[0,1]^{n}:\sum_{j=1}^{n}p_{j}=1\}$, where $p_{j}=\mathbb{P}(\{y_{j}\})$. Theorem 12 then states that regular proper scoring rules are all of the form
$S(\mathbf{p},y_{j})=H(\mathbf{p})+\langle h_{\mathbf{p}},\mathbf{e}_{j}-\mathbf{p}\rangle,$
[…]
Typically, the entropy $H$ is differentiable in the sense that for every $\mathbb{P}\in\mathcal{P}$, there exists a $\mathcal{P}$-integrable function $\nabla_{\mathbb{P}}H:\mathcal{Y}\to\overline{\mathbb{R}}$ such that
$\lim_{\alpha\to 0^{+}}\frac{1}{\alpha}\Big{[}H((1-\alpha)\mathbb{P}+\alpha\mathbb{Q})-H(\mathbb{P})\Big{]}=\langle\nabla_{\mathbb{P}}H,\mathbb{Q}\rangle$ (9)
for every $\mathbb{Q}\in\mathcal{P}$, and $\nabla_{\mathbb{P}}H$ is called the gradient of $H$. If $H$ is differentiable, then supergradients are unique and equal to the gradient $\nabla_{\mathbb{P}}H=h_{\mathbb{P}}$ (see Ekeland and Témam, 1999, Proposition 5.3). Thus, for differentiable entropies, there is a bijective correspondence between a scoring rule and its entropy and Equation (7) provides a generic recipe for constructing proper scoring rules out of differentiable concave functions on $\mathcal{P}$.
2.3 Geometric Properties
Generally, the divergence of a proper scoring rule does not correspond to a metric but behaves like a (squared) metric in many respects. Locally speaking, the divergence can be said to impose a manifold structure on the space of probability measures with the Hessian $\nabla_{\mathbb{P}}^{2}H$ of the entropy acting as the metric tensor by relating the “distance” $d(\mathbb{P},\mathbb{P}+\Delta\mathbb{P})$ between two infinitesimally close points $\mathbb{P},\mathbb{Q}\in\mathcal{P}$ to the difference $\Delta\mathbb{P}=\mathbb{Q}-\mathbb{P}$ between their “coordinates” $\mathbb{P}$ and $\mathbb{P}+\Delta\mathbb{P}$ as
$d(\mathbb{P},\mathbb{P}+\Delta\mathbb{P})\approx-\frac{1}{4}\langle\Delta\mathbb{P},[\nabla_{\mathbb{P}}^{2}H]\Delta\mathbb{P}\rangle=-\frac{1}{4}\int\nabla_{\mathbb{P}}^{2}H(x,x^{\prime})\,\mathrm{d}\Delta\mathbb{P}(x)\,\mathrm{d}\Delta\mathbb{P}(x^{\prime}),$ (10)
see Theorem 13 for a precise statement. Thus, the divergence acts like a (squared) metric locally between close enough points, with the metric depending on the points via the Hessian $\nabla_{\mathbb{P}}^{2}H$. This allows us to compare the local behavior of different scoring rules and has also been used to devise new scoring rules (see Bolin and Wallin, 2023).
Theorem 13.
Let $H:\mathcal{P}\to\overline{\mathbb{R}}$ be a twice-differentiable concave function and let $d:\mathcal{P}\times\mathcal{P}\to\overline{\mathbb{R}}$ denote the divergence of the corresponding scoring rule. Then
$d(\mathbb{P},\mathbb{Q})=-\frac{1}{2}\iint\left[\int_{0}^{1}t\nabla_{t\mathbb{P}+(1-t)\mathbb{Q}}^{2}H(x,x^{\prime})\,\mathrm{d}t\right]\,\mathrm{d}(\mathbb{Q}-\mathbb{P})(x)\,\mathrm{d}(\mathbb{Q}-\mathbb{P})(x^{\prime})$
for every $\mathbb{P},\mathbb{Q}\in\mathcal{P}$.
In other words, evaluating the divergence $d(\mathbb{P},\mathbb{Q})$ is a matter of integrating the Hessian over mixtures of $\mathbb{P}$ and $\mathbb{Q}$. This allows us to compare divergences of proper scoring rules to metrics which are more intuitive but less convenient in estimation.
Example 14.
Let $\mathbb{P},\mathbb{Q}$ be absolutely continuous distributions on $\mathbb{R}^{d}$ with Lebesgue densities $p$ and $q$, respectively. If $\mathbb{P}$ and $\mathbb{Q}$ are “close”, the divergence $d(\mathbb{P},\mathbb{Q})$ with respect to the logarithmic score, quadratic score (see Equation 16 in Section 3.3), and CRPS is approximately given by the right-hand side of Equation 10, which is, respectively,
$\frac{1}{4}\int(q(\mathbf{x})/p(\mathbf{x})-1)^{2}p(\textbf{x})\,\mathrm{d}\mathbf{x},\;\frac{1}{2}\int(q(\mathbf{x})-p(\mathbf{x}))^{2}\,\mathrm{d}\mathbf{x},\;-\frac{1}{8}\iint|x-y|(q(\mathbf{x})-p(\mathbf{x}))(q(\mathbf{y})-p(\mathbf{y}))\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mathbf{y}.$
The divergence also satisfies a generalization of the law of cosines (and the Pythagorean theorem). Indeed, the so-called three point property holds, which means that for $\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3}\in\mathcal{P}$,
$d(\mathbb{P}_{1},\mathbb{P}_{3})=d(\mathbb{P}_{1},\mathbb{P}_{2})+d(\mathbb{P}_{2},\mathbb{P}_{3})-\langle\nabla_{\mathbb{P}_{2}}H-\nabla_{\mathbb{P}_{1}}H,\mathbb{P}_{3}-\mathbb{P}_{2}\rangle.$
In particular, $d(\mathbb{P}_{1},\mathbb{P}_{3})=d(\mathbb{P}_{1},\mathbb{P}_{2})+d(\mathbb{P}_{2},\mathbb{P}_{3})$ if and only if $(\mathbb{P}_{1},\mathbb{P}_{2})$ and $(\mathbb{P}_{2},\mathbb{P}_{3})$ are geodesically orthogonal, that is $\langle\nabla_{\mathbb{P}_{2}}H-\nabla_{\mathbb{P}_{1}}H,\mathbb{P}_{3}-\mathbb{P}_{2}\rangle=0$. The geometry of proper scoring rules (like Bregman loss funcions) is said to be that of doubly flat manifolds, a class of Riemannian manifolds which captures the notion of “flatness” as exemplified by affine subspaces of Euclidean spaces (Dawid, 2007; Amari, 2016).
3 Families of Scoring Rules
3.1 Kernel scores
Kernel scores form a flexible and powerful class of proper scoring rules that generalize many familiar examples. They are based on reproducing kernel Hilbert space (RKHS) methods and offer a principled way to compare probability distributions using kernel-based embeddings. In what follows, we detail their construction and computational aspects.
3.1.1 Construction and computation
Kernel scores are a far-reaching generalization of the CRPS introduced in Example 3. They are derived from entropies of the form
$H(\mathbb{P})=\frac{1}{2}\iint h(x,x^{\prime})\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{P}(x^{\prime})$ (11)
for some conditionally negative definite kernel $h:\mathcal{Y}\times\mathcal{Y}\to[0,\infty)$, and have been popularized by Dawid (2007); Gneiting and Raftery (2007). A function $h:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is a conditionally negative definite kernel, if it is symmetric, that is, $h(x,y)=h(y,x)$ for $x,y\in\mathcal{Y}$, and for every $n\geq 1$, $x_{1},\dots,x_{n}\in\mathcal{Y}$ and $\alpha_{1},\dots,\alpha_{n}\in\mathbb{R}$ with $\sum_{j=1}^{n}\alpha_{j}=0$, it satisfies $\sum_{i,j=1}^{n}\alpha_{i}\alpha_{j}h(x_{i},x_{j})\leq 0$. Moreover, we assume that $h(x,y)\geq 0$ for $x,y\in\mathcal{Y}$. This does not lead to any loss of generality, see Remark 2 in the appendix.
Kernel scores are essentially squared distances between the kernel mean embeddings of the measures into some Hilbert space (Steinwart and Ziegel, 2021). They are closely related to maximum mean discrepancies (MMD), which were originally devised as test statistics for two-sample testing (Gretton et al., 2006). Furthermore, they are intimately connected to energy statistics (Székely and Rizzo, 2013). The relationship between energy statistics and MMDs is carefully studied in Sejdinovic et al. (2013), see also Lyons (2013).
For any conditionally negative definite kernel $h$, the entropy at Equation 11 is concave, and hence the induced scoring rule is proper. Strict propriety is a more delicate issue. Assume $h\geq 0$ and let $\mathcal{P}_{h}$ be the set of all probability measures such that $H(\mathbb{P})<\infty$. The kernel $h$ is strongly conditionally negative definite if $\iint h(x,y)\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(x)\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(y)=0$ implies $\mathbb{P}=\mathbb{Q}$ for any $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{h}$. Strongly conditionally negative definite kernels are closely related to positive definite characteristic and universal kernels, see Gretton et al. (2006); Sriperumbudur et al. (2011).
Theorem 15.
If $h:\mathcal{Y}\times\mathcal{Y}\to[0,\infty)$ is a (strongly) conditionally negative definite kernel, then the scoring rule $S:\mathcal{P}_{h}\times\mathcal{Y}\to\overline{\mathbb{R}}$ given by
$S(\mathbb{P},y)=\int h(y,x)\,\mathrm{d}\mathbb{P}(x)-\frac{1}{2}\iint h(x,x^{\prime})\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{P}(x^{\prime})$ (12)
is (strictly) proper.
A kernel score is a (strictly) proper scoring rule $S:\mathcal{P}_{h}\times\mathcal{Y}\to\overline{\mathbb{R}}$ of the form in Equation 12 for some (strongly) conditionally negative definite kernel $h:\mathcal{Y}\times\mathcal{Y}\to[0,\infty)$. The divergence $d$ of a kernel score is $d(\mathbb{P},\mathbb{Q})=-\frac{1}{2}\iint h(y,y^{\prime})\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(y)\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(y^{\prime}).$ We can write $S(\mathbb{P},y)=d(\mathbb{P},\delta_{y})+\tfrac{1}{2}h(y,y)$ since $\delta_{y}\in\mathcal{P}_{h}$ for $y\in\mathcal{Y}$. In fact, $(\mathbb{P},y)\mapsto d(\mathbb{P},y)$ is itself a proper scoring rule and is strongly equivalent to $S$. The intimate connection between conditionally negative definite kernels and the geometry of Hilbert spaces results in many interesting interpretations of the kernel score that are best stated in terms of its divergence.
Theorem 16 (Kernel Scores).
The function $S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ as in Equation 12 is a (strictly) proper scoring rule if and only if there exists a Hilbert space $\mathcal{H}$ and a subset $\{\psi_{x}\}_{x\in\mathcal{Y}}\subseteq\mathcal{H}$ such that the divergence $d$ of $S$ satisfies
$d(\mathbb{P},\mathbb{Q})=-\frac{1}{2}\iint\|\psi_{x}-\psi_{y}\|_{\mathcal{H}}^{2}\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(x)\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(y)=\left\|\int\psi_{x}\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(x)\right\|_{\mathcal{H}}^{2},$ (13)
(and that the mapping $\mathbb{P}\mapsto\int\psi_{x}\,\mathrm{d}\mathbb{P}(x)$ defined for $\mathbb{P}$ with $\int\|\psi_{x}\|^{2}\,\mathrm{d}\mathbb{P}(x)<\infty$ is injective).
We can thus construct a kernel score for a given outcome space $\mathcal{Y}$ as soon as we can construct a conditionally negative definite kernel on $\mathcal{Y}$ or a Hilbertian embedding of $\mathcal{Y}$. A simple approach is to choose $h=-k$ for some strictly positive definite kernel $k:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ which can be constructed in many ways (Williams and Rasmussen, 2006). For example, $k(\mathbf{x},\mathbf{y})=-\exp(-\|\mathbf{x}-\mathbf{y}\|^{p}/\lambda)$ corresponds to the Laplacian ($p=1$) and Gaussian ($p=2$) kernels on $\mathbb{R}^{d}$ leading to bounded kernel scores. Kernel scores implicitly correspond to different embeddings of the outcome space $\mathcal{Y}$ into a Hilbert space via Equation 13. For example, $(\mathbf{x},\mathbf{y})\mapsto\|\mathbf{x}-\mathbf{y}\|^{\beta/2}$ for $\beta\in(0,2)$ is actually a metric on $\mathcal{Y}=\mathbb{R}^{d}$. According to a well-known result of Schoenberg (1937), the corresponding metric space can be embedded into a Hilbert space, which implies the existence of an embedding $\{\psi_{\mathbf{x}}\}_{\mathbf{x}\in\mathcal{Y}}\subseteq\mathcal{H}$ such that $\|\psi_{\mathbf{x}}-\psi_{\mathbf{y}}\|_{\mathcal{H}}=\|\mathbf{x}-\mathbf{y}\|^{\beta/2}$ for $\mathbf{x},\mathbf{y}\in\mathcal{Y}$.
Example 17 (Energy Scores).
Let $\mathcal{Y}=\mathbb{R}^{d}$. For $\beta\in(0,2)$, the energy score $S_{\mathrm{Energy}}:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ given by
$S_{\mathrm{Energy}}(\mathbb{P},\mathbf{y})=\int\|\mathbf{x}-\mathbf{y}\|^{\beta}\,\mathrm{d}\mathbb{P}(\mathbf{x})-\frac{1}{2}\iint\|\mathbf{x}-\mathbf{x}^{\prime}\|^{\beta}\,\mathrm{d}\mathbb{P}(\mathbf{x})\mathbb{P}(\mathbf{x}^{\prime})$
[…]
Example 18 (Variogram Score).
Let $\mathcal{Y}=\mathbb{R}^{d}$. Scheuerer and Hamill (2015) propose the variogram score
$S(\mathbb{P},\mathbf{y})=\sum_{i,j=1}^{d}w_{ij}\left(|y_{i}-y_{j}|^{p}-\int|x_{i}-x_{j}|^{p}\,\mathrm{d}\mathbb{P}(\mathbf{x})\right)^{2}$
where $w_{ij}\geq 0$, $\mathbf{x}=(x_{i})_{i=1}^{d}$, $\mathbf{y}=(y_{i})_{i=1}^{d}$, and $p\in(0,\infty)$, defined on $\mathcal{P}=\{\mathbb{P}:\int|x_{i}|^{p}\,\mathrm{d}\mathbb{P}(\mathbf{x})<\infty,i=1,\dots,d\}$. It is a proper kernel score with conditionally negative definite kernel $h(x,y)=\sum_{i,j=1}^{d}w_{ij}(|x_{i}-x_{j}|^{p}-|y_{i}-y_{j}|^{p})^{2}$, see Allen et al. (2023). However, it is not strictly proper.
An advantageous feature of kernel scores is that they can be easily computed for discrete predictive distributions, or, generally, estimated by sampling from $\mathbb{P}$. Indeed, for an iid sample $X_{1},\dots,X_{m}$ from $\mathbb{P}$,
$\hat{S}(\mathbb{P},y)=\frac{1}{m}\sum_{i=1}^{m}h(X_{i},y)-\frac{1}{2m(m-1)}\sum_{i=1}^{m}\sum_{i^{\prime}:i^{\prime}\neq i}h(X_{i},X_{i^{\prime}})$ (14)
is an unbiased estimate of $\hat{S}(\mathbb{P},y)$. This provides an affordable alternative to numerical integration when a closed-form expression for $S(\mathbb{P},y)$ is not available.
3.1.2 Characterizations
The divergence of a proper scoring rule is generally not symmetric, although locally it behaves like a squared metric, as discussed in Section 2.3. Interestingly, the square root of the divergence of a kernel score is always a valid metric on $\mathcal{P}$. In fact, under mild conditions, these are the only proper scoring rules with this property.
Theorem 19.
Let $S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ be a regular proper scoring rule such that $\{\delta_{y}:y\in\mathcal{Y}\}\subseteq\mathcal{P}$ and the supergradient map $\mathbb{P}\mapsto h_{\mathbb{P}}$ is weakly continuous. Then, the divergence $d$ of $S$ is symmetric if and only if $S$ is strongly equivalent to a kernel score.
The definition of weak continuity of the supergradient map is given in the appendix. Theorem 19 need not be true when $\{\delta_{x}:x\in X\}\not\subseteq\mathcal{P}$ as exemplified by the quadratic score for Lebesgue densities on $\mathbb{R}^{d}$, see Section 3.3. It has divergence $d(\mathbb{P},\mathbb{Q})=\int(p(\mathbf{y})-q(\mathbf{y}))^{2}\,\mathrm{d}\mathbf{y}$ but is not a kernel score, see Remark 3 in the appendix.
We focus now on the outcome space $\mathcal{Y}=\mathbb{R}^{d}$. If we do not want to attach special importance to particular regions of $\mathcal{Y}=\mathbb{R}^{d}$, the scoring rule should be translation invariant, that is, $S(\mathbb{P},\mathbf{y})=S(\mathbb{P}_{\mathbf{h}},\mathbf{y}+\mathbf{h})$ for $\mathbf{y},\mathbf{h}\in\mathbb{R}^{d}$, where $\mathbb{P}_{\mathbf{h}}(A)=\mathbb{P}(A+\mathbf{h})$ for Borel sets $A\subseteq\mathbb{R}^{d}$. Such kernel scores correspond to kernels of the form $h(\mathbf{x},\mathbf{y})\equiv h(\mathbf{x}-\mathbf{y})$, $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, for some conditionally negative definite function $h:\mathbb{R}^{d}\to\mathbb{R}$. As a consequence of the Levy-Khinchin Theorem (see Sasvári, 2013, Theorem 3.12.12), they admit an elegant representation as the $L^{2}(\mu)$-distance of the characteristic functions, where $\mu$ is the spectral measure. This relates to the results of Székely et al. (2007). Similar results have independently been shown by Modeste and Dombry (2024) as the characterization of translation invariant MMDs on $\mathbb{R}^{d}$.
Theorem 20 (Spectral Representation of Translation Invariant Kernel Scores).
Let $S:\mathcal{P}\times\mathbb{R}^{d}\to\overline{\mathbb{R}}$ be a translation invariant kernel score. There exists a positive semi-definite matrix $\mathbf{B}\in\mathbb{R}^{d\times d}$, and a $\sigma$-finite measure $\mu$ on $\mathbb{R}^{d}$ with $\mu(\{\mathbf{0}\})=0$ and $\int_{\mathbb{R}^{d}}\min(1,\|\mathbf{u}\|^{2})\,\mathrm{d}\mu(\mathbf{u})<\infty$ such that
[…]
for every $\mathbf{y}\in\mathcal{Y}$, and $\mathbb{P},\mathbb{Q}\in\mathcal{P}$ with characteristic functions $f_{\mathbb{P}}$, $f_{\mathbb{Q}}$, and means $\mathbf{m}_{\mathbb{P}}$, $\mathbf{m}_{\mathbb{Q}}$, respectively. In particular, if the negative kernel is a positive definite function, then $\mu$ is a finite measure.
It follows that different translation invariant kernel scores place different weight on the different frequencies of the density. For example, the Gaussian and Laplacian kernel scores for $\lambda>0$, and the energy score for $\beta\in(0,2)$ in $\mathbb{R}^{d}$, correspond to $\mathbf{B}=\mathbf{0}$ and the spectral measures $\mu$ are given by
$\frac{\mathrm{d}\mu}{\mathrm{d}\mathbf{u}}\,\,\propto\,\,\exp(-\lambda\|\mathbf{u}\|^{2}),\quad\Big{(}1+\lambda^{2}\|\mathbf{u}\|^{2}\Big{)}^{-(d+1)/2},\mbox{ and }\|\mathbf{u}\|^{-(d+\beta)},$
respectively.
Further important notions of equivariance and invariance are homogeneity and isotropy. A proper scoring rule $S:\mathcal{P}\times\mathbb{R}^{d}\to\overline{\mathbb{R}}$ is said to be homogeneous of degree $\alpha$ if $S(\mathbb{P}_{c},c\mathbf{y})=c^{\alpha}S(\mathbb{P},\mathbf{y})$ for every $c>0$, $\mathbb{P}\in\mathcal{P}$, and $\mathbf{y}\in\mathbb{R}^{d}$ where $\mathbb{P}_{c}(A)=\mathbb{P}(c^{-1}A)$ for Borel sets $A\subseteq\mathbb{R}^{d}$. It is said to be isotropic or rotation-invariant if $S(\mathbb{P}_{\mathbf{U}},\mathbf{U}\mathbf{y})=S(\mathbb{P},\mathbf{y})$ for every rotation matrix $\mathbf{U}\in\mathrm{SO}(d)$, $\mathbb{P}\in\mathcal{P}$, and $\mathbf{y}\in\mathbb{R}^{d}$ where $\mathbb{P}_{\mathbf{U}}(A)=\mathbb{P}(\mathbf{U}^{\top}A)$ for Borel sets $A\subseteq\mathbb{R}^{d}$. A consequence of homogeneity, that is of great practical importance is that scaling the sample $\mathbf{y}$ by $c>0$ as $c\mathbf{y}$, scales the predictive distribution accordingly. Thus, using different units of measurement for the data (e.g. km/h and knots for wind speed) leads to the same optimal distribution in the corresponding units (in estimation contexts), or the same predictive performance assessment (in forecast evaluation contexts). The following result characterizes translation invariant kernel scores with these properties.
Theorem 21.
Let $S:\mathcal{P}\times\mathbb{R}^{d}\to\overline{\mathbb{R}}$ be a translation invariant kernel score with the matrix $\mathbf{B}$ and the spectral measure $\mu$ as in Theorem 20. Then there exists a $\sigma$-finite measure $\rho$ on (0,)(0$\infty$)( 0 , ∞ ) with $\int_{0}^{\infty}\min(1,r^{2})\,\mathrm{d}\rho(r)<\infty$ and probability measures $\{\nu_{r}:r\in(0,\infty)\}$ on the unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$ such that
$\,\mathrm{d}\mu(r\sigma)=\,\mathrm{d}\nu_{r}(\sigma)\,\mathrm{d}\rho(r)$
for $r\in(0,\infty)$ and $\sigma\in\mathbb{S}^{d-1}$.
  1. 1.
    If $S$ is homogeneous of degree $\alpha\in\mathbb{R}$, then either (a) $\alpha=2$ and $\mu$ is zero everywhere, or (b) $\alpha\in(0,2)$, $\mathbf{B}=\mathbf{0}$ and $\mu$ satisfies
    $\,\mathrm{d}\mu(r\sigma)=\alpha\mu(\mathbb{S}^{d-1}\times(0,1))\frac{\,\mathrm{d}\nu(\sigma)\,\mathrm{d}r}{r^{1+\alpha}}$
    for some probability measure $\nu$ on $\mathbb{S}^{d-1}$ and the Lebesgue measure $\mathrm{d}r$ on (0,)(0$\infty$)( 0 , ∞ ).
  2. 2.
    If $S$ is isotropic, then $\mathbf{B}=c\mathbf{I}$ for some $c\geq 0$ and $\mu$ satisfies
    $\,\mathrm{d}\mu(r\sigma)=\frac{\Gamma(d/2)}{2\pi^{d/2}}\,\,\mathrm{d}\sigma\,\,\mathrm{d}\rho(r)$
    where $\mathrm{d}\sigma$ is the (probability) Haar measure on $\mathbb{S}^{d-1}$.
In particular, the CRPS is the only 111-homogeneous translation invariant kernel score on $\mathbb{R}$ and the energy scores are the only homogeneous isotropic translation invariant kernel scores on $\mathbb{R}^{d}$ (up to positive multiplicative constants).
3.1.3 Continuous ranked probability score (CRPS)
The CRPS is arguably the most popular kernel score for univariate data. It was originally derived as the limit of discrete ranked probability scores and by integrating the Brier score, until it was recognized as a member of a much larger family of kernel scores (Staël von Holstein, 1977; Gneiting and Raftery, 2007). Mathematically, it is essentially the only 111-homogeneous translation invariant kernel score on $\mathbb{R}$ (see Theorem 21). In addition to historical reasons, its popularity is explained by a range of advantages that it offers.
The CRPS is an intuitive, interpretable, and convenient measure of forecasting accuracy. It is distance-sensitive, that is, it assigns better scores to forecasts, which put more probability closer to the observations. In contrast, the logarithmic score assigns harsh penalties regardless of the overall closeness of the forecast to the observations if the forecast density at any of the observations is too low. The CRPS is essentially the $L^{2}$ distance between the CDFs of the forecast and the observations (see Example 11), and as the integral of the quadratic (Brier) score for binary outcomes, it elegantly connects to the conceptually simpler point forecasting setting. Indeed, it reduces to the absolute error loss when point forecasts are issued. Moreover, homogeneity implies that scaling the observations scales the predictive distribution accordingly even when the CRPS is away from the minimum. Furthermore, the CRPS exists for all forecasts with a finite first moment, which means that the forecast measure need not necessarily have a density. It can be easily estimated by Equation 14 using samples drawn from the forecast, or explicitly computed if the forecast distribution has finite support. This is especially useful in applications such as ensemble forecasting in meteorology, where forecasts are routinely issued as discrete predictive distributions consisting of the empirical distributions of the ensemble of predictions.
The CRPS has been extended to circular random variables (Gneiting and Raftery, 2007). It has also spawned numerous other modifications, see Section 5.2.
Closed-form expressions for the CRPS have been obtained for many important parametric families such as the Gaussian and generalized extreme value (GEV) distributions (Taillardat et al., 2016). Closed-form expressions are also available for forecast CDFs which are linear splines (Gasthaus et al., 2019). In the absence of closed-form expressions, one can resort to estimating the score by sampling as in Equation 14 or numerical integration techniques such as kernel quadrature (Adachi et al., 2025). Directly implementing Equation 14 for the CRPS leads to an inefficient $\mathcal{O}(n^{2})$ time computation where $n$ is the number of samples drawn from the forecast. This can be improved to almost linear $\mathcal{O}(n\log n)$ time using an algebraically equivalent expression (Kelen et al., 2025):
$\mathrm{CRPS}(\hat{\mathbb{P}},y)=\frac{2}{n(n-1)}\sum_{j=1}^{n}(X_{(j)}-y)\left((n-1)\mathbf{1}_{\{y<X_{(j)}\}}-j+1\right)$
where ^$\hat{\mathbb{P}}=\frac{1}{n}\sum_{j=1}^{n}\delta_{X_{j}}$.
The CRPS plays a central role in forecast verification in meteorology. Moreover, with the widespread adoption of weather-related renewable energy sources such as solar and wind, accurate probabilistic forecasts for power, load and price of energy are necessary due to their highly variable nature (Zhang et al., 2014; Hong and Fan, 2016; Weron, 2014). In these applications, the CRPS is widely used as an evaluation metric and increasingly also for estimation. Furthermore, the CRPS has also been used for parametric (Gneiting et al., 2005) and nonparametric (Henzi et al., 2021) regression in statistics, and for supervised learning of distributions in machine learning (Gasthaus et al., 2019). We discuss the role and use of proper scoring rules in estimation further in Section 5.
Pic et al. (2023) derived the minimax rate of convergence for distributional regression with the CRPS and showed that it is achieved by using $k$-nearest neighbors ($k$-NN) or kernel smoothing. Surprisingly, this rate is of the same order as the minimax optimal rate for point prediction. Theoretical guarantees for estimation, model selection and convex aggregation have been obtained by Clement and Zaoui (2024).
3.2 Local scoring rules
A local scoring rule is a scoring rule whose value $S(\mathbb{P},x)$ depends on $\mathbb{P}$ only through the values of its density on an arbitrarily small neighborhood around $x$. Let $\mathcal{Y}\subseteq\mathbb{R}^{d}$ be open and let $\mathcal{P}_{\mathrm{ac}}^{k}$ denote the set of probability measures on $\mathcal{Y}$ with Lebesgue densities, which are $k$-times differentiable.
Definition 22.
A local scoring rule of order $k$ is a proper scoring rule $S:\mathcal{P}_{\mathrm{ac}}^{k}\times\mathcal{Y}\to\overline{\mathbb{R}}$ of the form
$S(\mathbb{P},\mathbf{y})=s(\mathbf{y},p(\mathbf{y}),\nabla_{\mathbf{y}}\,p(\mathbf{y}),\dots,\nabla_{\mathbf{y}}^{k}\,p(\mathbf{y}))$
where $s:\mathcal{Y}\times\mathbb{R}\times\cdots\times\mathbb{R}^{d^{k}}\to\mathbb{R}$, $p$ is the density of $\mathbb{P}$ and $\nabla_{\mathbf{y}}^{j}\,p(\mathbf{y})$ denotes the $j$th derivative of $p$.
Arguably, the most popular and important of all proper scoring rules is the logarithmic score given by $S_{\mathrm{log}}(\mathbb{P},\mathbf{y})=-\log p(\mathbf{y})$, compare Examples 5 and 9. It is a local scoring rule of order 0, and remarkably, the only differentiable scoring rule of this kind (Bernardo, 1979). Empirical risk minimization for the logarithmic score is equivalent to maximum likelihood estimation which is the foundation of classical parametric inference. The unparalleled historical success of maximum likelihood is in great part because it furnishes closed-form estimates for many important parametric families of distributions and is asymptotically optimal under rather mild regularity conditions.
Remark 23.
Theorem 12 can be used to derive a simple proof that the logarithmic score is the only differentiable local scoring rule of order 0. Indeed, let $S(\mathbb{P},\mathbf{y})=s(\mathbf{y},p(\mathbf{y}))$ be such a scoring function. The function $s(\mathbf{y},z)$ is differentiable in $z$, and its entropy is $H(\mathbb{P})=\int s(\mathbf{x},p(\mathbf{x}))p(\mathbf{x})\,\mathrm{d}\mathbf{x}$. By Equation (7) and using Equation (9), we must have
$S(\mathbb{P},\mathbf{y})=s(\mathbf{y},p(\mathbf{y}))+p(\mathbf{y})\frac{\partial}{\partial z}s(\mathbf{y},p(\mathbf{y}))-\int\frac{\partial}{\partial z}s(\mathbf{x},p(\mathbf{x}))[p(\mathbf{x})]^{2}\,\mathrm{d}\mathbf{x},$
which is of the form $s(\mathbf{y},p(\mathbf{y}))$ if and only if the integral is a constant, or equivalently, the integrand is a constant multiple of $p(\mathbf{y})$. It follows that $s(\mathbf{y},p(\mathbf{y}))=c\log p(\mathbf{y})+d$ for some constants $c,d\in\mathbb{R}$.
Even though the logarithmic score is the preferred approach to fitting standard parametric models, it is frequently impractical for fitting non-standard ones because the normalizing constants of such families are often not available in closed form and estimating or approximating them is computationally infeasible. To circumvent this problem, Hyvärinen (2005) proposed to fit models using score matching, which is to minimize the expected squared distance $\mathbb{E}_{\mathbf{X}\sim p_{\theta}}[\|\nabla_{\mathbf{x}}\log p_{\theta}(\mathbf{X})-\nabla_{\mathbf{x}}\log q(\mathbf{X})\|^{2}]$ between the “scores” $\nabla_{\mathbf{x}}\log p_{\theta}(\mathbf{x})$ and $\nabla_{\mathbf{x}}\log q(\mathbf{x})$, where $q$ is the true density. Empirically, this corresponds to minimizing the Hyvärinen score given by
$S_{\mathrm{Hyv\ddot{a}rinen}}(\mathbb{P},\mathbf{y})=\Delta_{\mathbf{y}}\log p(\mathbf{y})+\tfrac{1}{2}\|\nabla_{\mathbf{y}}\log p(\mathbf{y})\|^{2},$
where $p$ is the density of $\mathbb{P}$, see Examples 6 and 10. It is a local scoring rule of order 222. The value of the Hyvärinen score does not change if we replace $p$ by $c\cdot p$ for some constant $c>0$, which implies that the score can be evaluated even with non-normalized $p$.
The Hyvärinnen score has also been extended to distributions on $\{-1,1\}^{d}$, $\mathbb{R}_{+}^{n}$ (Hyvärinen, 2007), spheres (Takasu et al., 2018) and general Riemannian manifolds (Dawid and Lauritzen, 2005; Mardia et al., 2016). Due to its applicability to non-normalized models, score matching enjoys significant popularity in the machine learning community and has, in fact, played an important role in recent advances in generative modeling, see Section 5.1.1.
Forbes and Lauritzen (2015) establish the asymptotic normality of parameter estimation using the Hyvärinnen score and Koehler et al. (2022) relate the asymptotic efficiency of the estimator to the isoperimetric constant of the distribution. Ghosh et al. (2025) establish minimax rate optimality (up to log factors) of denoising.
A fairly general characterization of local scoring rules of arbitrary order has been obtained by Parry et al. (2012) using a calculus of variations approach. In particular, there are no local scoring rules of odd orders, and all local scoring rules of even order 2$\geq$ 2≥ 2 can be computed without knowledge of the normalization constant of the density like in the case of the Hyvärinen score. A more explicit characterization for the particular case of order 222 was derived by Ehm and Gneiting (2012) using an approach based on Theorem 12. For example, Ehm and Gneiting (2012, Example 5.3) introduce the log-cosh score
$S_{\text{log-cosh}}(\mathbb{P},y)=-\log\cosh z_{1}+z_{1}\tanh z_{1}+z_{2}(1-\tanh^{2}z_{1}),$
with $z_{1}=(\mathrm{d}/\mathrm{d}y)\log p(y)=p^{\prime}(y)/p(y)$ and $z_{2}=(\mathrm{d}^{2}/\mathrm{d}y^{2})\log p(y)=p^{\prime\prime}(y)/p(y)-(p^{\prime}(y)/p(y))^{2}$. It is strictly proper for distributions $\mathbb{P}$ on $\mathbb{R}$ with densities $p$ that satisfy some regularity conditions (Ehm and Gneiting, 2012, Definition 2.3). Parry (2016) studied local scoring rules in the multivariate setting. Specifically, he introduced and characterized the so-called extensive scoring rules, which are multivariate scoring rules that decompose into sums of univariate scoring rules for individual components. Dawid et al. (2012) characterized scoring rules for general discrete outcome spaces which are local in that they depend only on the values of the probability mass function on a neighborhood of $x$ in a given undirected graph.
3.3 $f$-Scores, $gf$-scores and generalized kernel scores
Let $\mu$ be a $\sigma$-finite measure on $\mathcal{Y}$ and $\mathcal{P}_{\mathrm{ac}}$ the set of measures which are absolutely continuous with respect to $\mu$. For example, if $\mathcal{Y}=\mathbb{R}^{d}$ and $\mu$ is the Lebesgue measure, we consider absolutely continuous distributions on $\mathbb{R}^{d}$; if $\mathcal{Y}$ is finite and $\mu$ is the counting measure, we have categorical random variables.
Given a concave and differentiable function $f:[0,\infty)\to\mathbb{R}$, we can define a concave entropy function $H_{f}:\mathcal{P}\to\overline{\mathbb{R}}$ as
$H_{f}(\mathbb{P})=\int f(p(x))\,\mathrm{d}\mu(x),$ (15)
where $p$ denotes the density of $\mathbb{P}\in\mathcal{P}\subseteq\mathcal{P}_{\mathrm{ac}}$ such that $H_{f}(\mathbb{P})<\infty$ for $\mathbb{P}\in\mathcal{P}$. Using Theorem 12 and Equation (9), it follows that the corresponding proper scoring rule $S_{f}:\mathcal{X}\times\mathcal{P}\to\overline{\mathbb{R}}$ is given by
$S_{f}(\mathbb{P},y)=f^{\prime}(p(y))+H_{f}(\mathbb{P})-\int f^{\prime}(p(x))p(x)\,\mathrm{d}\mu(y).$
Such scoring rules were introduced by Grünwald and Dawid (2004) as separable Bregman scores, although we refer to them as $f$-scores to avoid confusion with Bregman loss functions and divergences. Pérez (1967) introduced the entropies in Equation 15 as generalized $f$-entropies. However, they are different from the $f$-entropies of Csiszár (1972). $f$-scores are the only proper scoring rules of the form $S(\mathbb{P},y)=g(p(y))-l(\mathbb{P})$ (Dawid, 2007).
Many important scores are instances of this family. The logarithmic score arises for $f(u)=-u\log u$, compare Example 5 and Section 3.2. The concave function $f(u)=-u^{2}$ leads to the well-known Brier or quadratic score
$S_{\mathrm{Brier}}(\mathbb{P},y)=-2p(y)+\int p(x)^{2}\,\mathrm{d}\mu(x),$ (16)
compare Example 1, where it is stated in an equivalent form for binary outcomes.
An extension of $f$-scores are $gf$-scores with entropy functions of the form
$H_{gf}(\mathbb{P})=g\Big{(}\int f(p(x))\,\mathrm{d}\mu(x)\Big{)}=g(H_{f}(\mathbb{P})),$ (17)
where $f$ and $g$ are such that $H$ is concave. For example, $f$ could be concave and $g$ concave and increasing, or, $f$ convex and $g$ concave and decreasing. If $f$ and $g$ are differentiable, applying Theorem 12 and Equation (9) yields
$S_{gf}(\mathbb{P},y)=g^{\prime}(H_{f}(\mathbb{P}))f^{\prime}(p(y))+g(H_{f}(\mathbb{P}))-g^{\prime}(H_{f}(\mathbb{P}))\int f^{\prime}(p(x))p(x)\,\mathrm{d}\mu(x).$
The spherical (Roby, 1964) and pseudospherical scoring rules (Good, 1971) given by
$S_{\mathrm{Spherical}}(\mathbb{P},y)=-\frac{p(y)}{\left[\int p(x)^{2}\,\mathrm{d}\mu(x)\right]^{1/2}}\quad\text{and}\quad S_{\mathrm{pSpherical}}(\mathbb{P},y)=-\frac{p(y)^{\alpha-1}}{\left[\int p(y)^{\alpha}\,\mathrm{d}\mu(x)\right]^{1-1/\alpha}},$
respectively are instances of $gf$-scores with $f(u)=u^{\alpha}$ and $g(u)=-u^{1/\alpha}$ for $\alpha=2$ and $\alpha>1$, respectively. In this case, strict concavity of the entropy follows from the Minkowski inequality.
The logarithmic, quadratic and spherical scores have been axiomatically characterized by Shuford et al. (1966), Selten (1998) and Jose (2009), respectively. Recently, Shao et al. (2024) explored their application to training large language models (LLMs).
Another twist on the idea is to replace the integral in Equation 17 with a different entropy, for example, the entropy of a kernel score as at Equation 11, that is, considering entropies of the form
$H(\mathbb{P})=g\left(\frac{1}{2}\iint h(x,x^{\prime})\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{P}(x^{\prime})\right)=g(H_{[h]}(\mathbb{P}))$
for a conditionally negative definite kernel $h$ and a differentiable increasing concave function $g:[0,\infty)\to\mathbb{R}$. This leads to the generalized kernel scores given by
$S_{g[h]}(\mathbb{P},y)=g^{\prime}(H_{[h]}(\mathbb{P}))\int h(x,y)\,\mathrm{d}\mathbb{P}(x)+g(H_{[h]}(\mathbb{P}))-g^{\prime}(H_{[h]}(\mathbb{P}))H_{[h]}(\mathbb{P}),$
introduced by Bolin and Wallin (2023).
3.4 Proper scoring rules induced by elicitable functionals
Proper scoring rules can also be constructed by reducing the probability measure $\mathbb{P}$ first to some summary measure, a functional $T:\mathcal{P}\to\mathcal{Z}$ and then comparing distributions by comparing their summaries. Such scoring rules are typically proper but not strictly proper.
A functional $T:\mathcal{P}\to\mathcal{Z}$ is called elicitable if there exists a scoring function $s:\mathcal{Z}\times\mathcal{Y}\to\overline{\mathbb{R}}$ such that
$\mathbb{E}s(T(\mathbb{P}),Y)\leq\mathbb{E}s(z,Y)$
for all $z\in\mathcal{Z}$, $\mathbb{P}\in\mathcal{P}$, where $Y\sim\mathbb{P}$, subject to the existence of expectations, with equality if and only if $z=T(\mathbb{P})$. Such a scoring function $s$ is called strictly consistent for the functional $T$. Relevant functionals that are elicitable are the mean, quantiles, expectiles, and ratios of expectations. Some important functionals like the variance are not elicitable. There is a large literature on elicitability with possible starting points being Lambert et al. (2008); Gneiting (2011); Fissler and Ziegel (2016).
Given a strictly consistent scoring function $s$ of an elicitable functional $T$, the scoring rule $S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ given by
$S(\mathbb{P},y)=s(T(\mathbb{P}),y)$
is proper, see Gneiting (2011, Theorem 3). For example, since the squared error $s(x,y)=(x-y)^{2}$ is strictly consistent for the mean, the scoring function $(\mathbb{P},y)\mapsto(\int x\,\mathrm{d}\mathbb{P}(x)-y)^{2}$ is proper but it is clearly not strictly proper unless the class $\mathcal{P}$ is a location family.
Example 24 (Dawid-Sebastiani Scores).
The pair of expectation and covariance matrix of multivariate distributions is elicitable. The resulting proper scoring rules have been introduced by Dawid and Sebastiani (1999). An interesting member of this class is given by
$S_{\mathrm{DS}}(\mathbb{P},\mathbf{y})=\log\det\Sigma_{\mathbb{P}}+(\mathbf{y}-m_{\mathbb{P}})^{\top}\Sigma_{\mathbb{P}}^{-1}(\mathbf{y}-m_{\mathbb{P}}),$
where $\mathbf{y}\in\mathbb{R}^{d}$ and $m_{\mathbb{P}}=\mathbb{E}[\mathbf{X}]$ and $\Sigma_{\mathbb{P}}=\mathbb{E}[(\mathbf{X}-m_{\mathbb{P}})(\mathbf{X}-m_{\mathbb{P}})^{\top}]$ are the mean and covariance of $\mathbf{X}\sim\mathbb{P}$. It is assumed that the covariance matrix is non-singular. The proper scoring rule $S_{\mathrm{DS}}$ is equivalent to the negative log-likelihood of a multivariate Gaussian distribution. As in the univariate case, it is a proper scoring rule for all distributions with non-singular covariance matrix.
4 Scoring Rule Decompositions
Proper scoring rules can be decomposed into interpretable components that provide insight concerning the performance of a given method. In the context of estimation, it is natural to consider an analogue of a (conditional) bias-variance decomposition, whereas in forecast evaluation, proper scoring rules are typically decomposed into components quantifying miscalibration, discrimination ability, and uncertainty.
[…]
$\mathbb{E}[d(\smash{\mathbb{P}_{\hat{\theta},X}},\mathbb{P}_{Y|X})]=\underbrace{\mathbb{E}[d(\smash{\mathbb{P}_{\hat{\theta},X}},\bar{\mathbb{P}}^{\phantom{\ast}}_{X})]}_{\mathrm{variance}}+\underbrace{\mathbb{E}[d(\bar{\mathbb{P}}^{\phantom{\ast}}_{X},\mathbb{P}_{Y|X})]}_{\mathrm{bias}}$
where ¯$\bar{\mathbb{P}}_{x}=\operatorname*{arg\,min}_{\mathbb{P}}\mathbb{E}[d(\smash{\mathbb{P}_{\hat{\theta},x}},\mathbb{P})]$ and $\mathbb{P}_{Y|X=x}=\operatorname*{arg\,min}_{\mathbb{P}}\mathbb{E}[S(\mathbb{P},Y)\,|\,X=x]$ is the best predictor, assuming the two exist. If we think of ¯$\smash{\bar{\mathbb{P}}_{x}}$ as the “expectation” of the predictor $\smash{\mathbb{P}_{\hat{\theta},x}}$ over $\mathcal{D}$, then just like in the original bias-variance decomposition, the variance term measures how much the estimator $\smash{\mathbb{P}_{\hat{\theta},x}}$ varies around ¯$\smash{\bar{\mathbb{P}}_{x}}$ and the bias term measures the distance between ¯$\smash{\bar{\mathbb{P}}_{x}}$ and the best predictor $\mathbb{P}_{Y|X=x}$. Indeed, the two decompositions coincide for the Brier score. While ¯$\bar{\mathbb{P}}_{x}$ corresponds to a left Bregman projection, commonly studied in optimization, the optimal predictor $\mathbb{P}_{Y|X=x}$ is a right Bregman projection, which is of greater importance in statistics (Bauschke et al., 2018).
In meteorology, a different kind of decomposition is used for assessing the overall performance of probabilistic forecasts, see Bröcker (2009); Kull and Flach (2015); Arnold et al. (2024) and references therein. The expected proper scoring rule is decomposed into terms corresponding to discrimination ability (DSC, or resolution), miscalibration (MCB, or reliability) and uncertainty (UNC). Let $\hat{\mathbb{P}}$ denote a (random) probabilistic forecast of $Y$. Then we can write
$\mathbb{E}[S(\hat{\mathbb{P}},Y)]=\underbrace{\,\mathbb{E}[d(\hat{\mathbb{P}},\mathbb{P}_{\smash{\raisebox{-3.01389pt}{$\hat{}\,$}}})]\,}_{\mathrm{MCB}}-\underbrace{\,\mathbb{E}[d(\bar{\mathbb{P}},\mathbb{P}_{\smash{\raisebox{-3.01389pt}{$\hat{}\,$}}})]\,}_{\mathrm{DSC}}+\underbrace{\,H(\bar{\mathbb{P}}_{\phantom{\smash{\raisebox{-2.71246pt}{$\hat{}\,$}}}})}_{\mathrm{UNC}\,}$
where $\mathbb{P}_{\smash{\raisebox{-3.61664pt}{$\hat{}\,$}}}$ denotes the conditional law of the outcome $Y$ given the forecast $\hat{\mathbb{P}}$ and $\bar{\mathbb{P}}$ is the marginal law of $Y$, often called the climatological forecast in meteorology. The miscalibration (MCB) term measures how different the forecast $\hat{\mathbb{P}}$ is from the optimal calibrated forecast $\mathbb{P}_{\smash{\raisebox{-3.61664pt}{$\hat{},$}}}$ with the same information, which is the conditional law of $Y$ given $\hat{\mathbb{P}}$. Discrimination ability (DSC) quantifies how much more discriminative the optimal calibrated forecast $\mathbb{P}_{\smash{\raisebox{-3.61664pt}{$\hat{}\,$}}}$ for the given information is in comparison to only marginal information about $Y$. Finally, the uncertainty term (UNC) informs about the general difficulty of the forecasting problem.
5 Applications
In this section, we briefly review the many ways in which proper scoring rules can be used to address computational and statistical considerations in estimation and forecast evaluation contexts.
5.1 Computational Tractability
There are often more convenient and effective ways to model a probability distribution than as a member of a standard parametric family. Estimation of such models using the logarithmic score in the form of maximum likelihood can, however, present an often insurmountable computational burden. Proper scoring rules can provide computationally viable alternatives.
5.1.1 Non-normalized Densities
Although the logarithmic score yields closed-form expressions for many standard parametric families, it is often intractable for more complicated ones. This occurs frequently in statistics (see for example de Fondeville and Davison, 2018) and especially in some modern machine learning applications such as generative diffusion models (Song and Ermon, 2019), where we need to model probability densities more flexibly, say as $p_{\theta}(\mathbf{x})\propto\exp\eta_{\theta}(\mathbf{x})$, where $\eta_{\theta}$ could be a neural network. Estimating $\theta$ via maximum likelihood for such complicated models would require evaluating the normalization constant $c_{\theta}=\int\exp\eta_{\theta}(\mathbf{x})\,\mathrm{d}\mathbf{x}$, which is prohibitively expensive even for moderately high dimensions. The Hyvärinnen score allows for a much more tractable estimation of such models by circumventing the need to evaluate the normalization constant $c_{\theta}$ and reducing the problem to minimizing the empirical risk corresponding to
$\Delta_{\mathbf{x}}\eta_{\theta}(\mathbf{x})+\frac{1}{2}\|\nabla_{\mathbf{x}}\eta_{\theta}(\mathbf{x})\|^{2}$
which depends only on $\eta_{\theta}$. In this way, it complements methods like Langevin dynamics, which allow sampling from $p_{\theta}$ using only $\eta_{\theta}$. This makes it particularly convenient for many statistical applications where normalization can be an issue such as estimation of high-dimensional graphical models (Lin et al., 2016), causal discovery (Rolland et al., 2022) and estimation of point processes (Cao et al., 2024). In Bayesian statistics, the Hyvärinen score has been used for likelihood-free inference (Pacchiardi and Dutta, 2022), model selection (Dawid and Musio, 2015), the derivation of objective priors (Leisen et al., 2020), model comparison (Shao et al., 2019) and sampling from Markov chain Monte Carlo and variational inference approaches (Zhang et al., 2018; Modi et al., 2023). The recent emergence of diffusion models as a powerful approach to generative modeling of images, audio, video and text, is also partly due to advancements in the implementation of score matching (see for example Dhariwal and Nichol, 2021).
The pseudo-spherical scoring rules also enable estimation with non-normalized densities due to their homogeneity and for this reason, they have been applied to estimating distributions on discrete spaces (Takenouchi and Kanamori, 2015) and training energy-based models (Yu et al., 2021).
5.1.2 Transformation Models
In many settings in statistics and machine learning, such as that of structural equation models (Dimitruk et al., 2007) or implicit generative models (Mohamed and Lakshminarayanan, 2017), we model probability distributions $\{\mathbb{P}_{\theta}\}_{\theta\in\Theta}$ (or conditional distributions $\{\mathbb{P}_{\theta,\mathbf{x}}\}_{\theta\in\Theta}$) as the laws of parameterized transformations $g_{\theta}(\mathbf{N})$ (or conditionally as $g_{\theta}(\mathbf{x},\mathbf{N})$) for $\theta\in\Theta$, where $\mathbf{N}$ is some standard random vector such as a multivariate standard Gaussian, instead of a parametric family of densities $\{p_{\theta}:\theta\in\Theta\}$. Fitting such probability distributions by estimating $\theta$ using maximum likelihood via a change of variables would require potentially expensive Jacobian computations and is difficult to implement if the transformations $g_{\theta}$ are not bijective or do not have a closed-form inverse. The original approach to generative adversarial networks (GANs) achieved this via an innovative min-max formulation which however complicates training (Goodfellow et al., 2014). Kernel scores have proven to be a viable alternative approach to training such models since they can be estimated directly by sampling from $\mathbf{N}$ as
$\hat{S}(\mathbb{P}_{\theta,\mathbf{x}},y)=\frac{1}{m}\sum_{i=1}^{m}h(g_{\theta}(\mathbf{x},\mathbf{N}_{i}),y)-\frac{1}{2m(m-1)}\sum_{i=1}^{m}\sum_{i^{\prime}:i^{\prime}\neq i}h(g_{\theta}(\mathbf{x},\mathbf{N}_{i}),g_{\theta}(\mathbf{x},\mathbf{N}_{i^{\prime}}))$
in the conditional case, for example, where $\{\mathbf{N}_{i}\}_{i=1}^{m}$ are independent realizations of $\mathbf{N}$, without requiring the density or an unstable adversarial formulation like GANs (Bouchacourt et al., 2016; Gritsenko et al., 2020).
Consequently, they have been found to be very useful for probabilistic forecasting (Pacchiardi et al., 2024a), estimating conditional transformation models (Hothorn et al., 2014), postprocessing of numerical weather prediction models (Chen et al., 2024), Bayesian generalized likelihood-free inference (Pacchiardi et al., 2024b), extrapolating neural networks (Shen and Meinshausen, 2024), uncertainty quantification of dynamical systems (Bülte et al., 2024), training neural SDEs (Issa et al., 2023), representation learning (Vahidi et al., 2024) and even estimating generative diffusion models (Bortoli et al., 2025). Although Gneiting et al. (2005) was the first to apply kernel score minimization for estimating a parametric family, the earliest instance of fitting conditional transformation models appears to be Hothorn et al. (2014) and Bouchacourt et al. (2016) (for neural networks). Non-conditional transformation models based on neural networks appear to have been first estimated using kernel scores by Dziugaite et al. (2015) and Li et al. (2015).
5.2 Statistical considerations
In the literature on forecast evaluation, many proper scoring rules have been constructed with the aim of detecting specific forecast deficiencies, or for targeting specific types of outcomes such as high impact events. Although these scoring rules have been shown to be useful in evaluation, they have not been widely used in estimation contexts.
5.2.1 Sensitivity and Robustness
Many scoring rules have been proposed in order to be sensitive or robust to particular statistical properties of the observations. For example, Scheuerer and Hamill (2015) proposed the variogram score in order to remedy the relative insensitivity of the energy scores to the correlation structure of multivariate distributions leading to poor discrimination ability (Pinson and Tastu, 2013). In contrast, Bolin and Wallin (2023) considered a setting where observations are drawn from different distributions, where they show that common scoring rules such as the CRPS overweigh observations coming from more dispersed distributions. Motivated by local geometry considerations (see Section 2.3), they proposed the scaled CRPS to correct for this relative oversensitivity to observation variance which can be understood as a nuisance parameter. Another modification known as the scaled weighted CRPS has been proposed by Olafsdottir et al. (2024) in this setting to account for extremes.
5.2.2 Weighted Scoring Rules
In many statistical applications such as risk management (Holzmann and Klar, 2017) and high impact weather prediction (Allen et al., 2023), certain outcomes are of considerably greater consequence to the user than others and naturally, we want the scoring rule to weigh forecast performance for such outcomes accordingly. To achieve this, Allen et al. (2023) propose using kernel scores with kernels modified using threshold-weighting or vertical-rescaling, that is, $h^{\prime}(x,y)=h(v(x),v(y))$ and $h^{\prime}(x,y)=h(x,y)w(x)w(y)$, respectively, by using a chaining function $v:\mathcal{Y}\to\mathcal{Y}$ or weighting function $w:\mathcal{Y}\to\mathbb{R}$. By choosing $h$ to be a translation invariant kernel, we can control the weight exclusively through the chaining and weighting functions. This generalizes the threshold-weighted CRPS introduced by Gneiting and Ranjan (2011), which also introduced a quantile-weighted version of the CRPS. Towards the same end, Diks et al. (2011) propose weighted versions of the logarithmic score based on which Holzmann and Klar (2017) propose a different approach to modifying scoring rules, which generally does not lead to kernel scores even if the original score was a kernel score.
For binary outcomes, Buja et al. (2005) argued that scoring rules should be tailored to the relative cost implications of different types of errors. In a decision-theoretic context, Johnstone et al. (2011) proposed tailored scoring rules which are proper scoring rules adapted to the utility function of the decision maker.
6 Limitations and Extensions
Proper scoring rules are versatile and important tools for estimation and forecast evaluation. However, they also have limitations. Proper scoring rules cannot assess tail properties of distributions. For example, under weak conditions on the scoring rule $S$, Brehmer and Strokorb (2019) demonstrate that for a given distribution $\mathbb{P}$ on $\mathbb{R}$, one can always find a distribution $\mathbb{Q}$ such that the divergence $d(\mathbb{Q},\mathbb{P})=S(\mathbb{Q},\mathbb{P})-S(\mathbb{P},\mathbb{P})$ is arbitrarily small but $\mathbb{P}$ and $\mathbb{Q}$ are not tail equivalent. This negative result has further been strengthened by Allen et al. (2024).
In survival analysis, the classical approach to inference based on maximum likelihood (i.e. the logarithmic score) is complicated by the fact that observations are censored because this often renders the likelihood intractable. Consequently, many approaches based on alternative scoring rules such as the survival-CRPS (Avati et al., 2020) and integrated Brier score (Graf et al., 1999) have been devised over the years, see also Dawid and Musio (2014). Unfortunately, these have been found to be improper by Rindt et al. (2022), implying that they are unfit for evaluation or estimation. Rindt et al. (2022) establish the propriety of the classical right-censored log-likelihood and propose a tractable approach to survival analysis based on it. Moreover, Yanagisawa (2023) proposed proper extensions of the Brier and ranked probability scores to survival analysis.
Hanson (2003) has proposed an interesting extension of proper scoring rules called market scoring rules for improving prediction markets. The basic premise of a prediction market is as follows: for a binary event of interest, say if it will rain on a given day, one can float a security in the prediction market which pays one dollar if (and when) the event takes place and nothing if it does not. The price of such a security can then be interpreted as the probability of the event (as forecasted by the market). However, obtaining the market forecast for a continuous quantity, such as the amount of rainfall on a given day, or a joint forecast for a multitude of related events, would require a very large number of such binary securities. Consequently, each security will barely have any buyers or sellers and the market will not function. This is known as the thin market problem. Hanson’s solution is to use instead a single security based on a proper scoring rule $S$ on the possible probability distributions of the outcomes and an automated market maker (AMM) which functions as follows: if a participant disagrees with the AMM forecast $\mathbb{P}_{t}$ at time $t$, he can replace it with his own $\mathbb{P}_{t+1}$ and be paid [^2] We use scores in their positive orientation here. $S(\mathbb{P}_{t+1},y)-S(\mathbb{P}_{t},y)$ if the random quantity realized is $y$. The idea has received renewed interest due to recent developments concerning cryptocurrencies. Interesting further developments include Hanson (2007), Abramowicz (2007) and Berg and Proebsting (2009).
Acknowledgements
The authors would like to thank an anonymous reviewer, Sam Allen, Philip Dawid, Ritabrata Dutta, Sebastian Engelke, Rafael Frongillo, Tilmann Gneiting, Daniel Hug, Pierre Jacob, Alexander Jordan, Kristof Kraus, Andrea Nava, Romain Pic, Mathias Trabs, and Xinwei Shen for comments, inspiring and fruitful discussions, and useful pointers to the literature.
Appendix
Appendix A Proofs for Section 2
Proof of Theorem 8.
If $S$ is a (strictly) proper scoring rule, then, by Fubini’s theorem,
$\mathbb{E}[S(\mathbb{P}^{X},Y)]$ $=\int_{\mathcal{X}}\int_{\mathcal{Y}}S(\mathbb{P}^{x},y)\,\mathrm{d}\mathbb{P}_{Y|X=x}(y)\,\mathrm{d}\mathbb{P}_{X}(x)$
$=\int_{\mathcal{X}}S(\mathbb{P}^{x},\mathbb{P}_{Y|X=x})\,\mathrm{d}\mathbb{P}_{X}(x)$
is minimized if (and only if) $\mathbb{P}^{x}=\mathbb{P}_{Y|X=x}$ for $x\in\mathcal{X}$ $\mathbb{P}_{X}$-almost surely. For the converse, we only need to take $X$ constant. ∎
Remark 1 (Derivation of Hyvärinen score entropy).
Clearly,
$H(\mathbb{P})$ $=\int\left[\Delta_{\mathbf{y}}\log p(\mathbf{y})+\frac{1}{2}\|\nabla_{\mathbf{y}}\log p(\mathbf{y})\|^{2}\right]p(\mathbf{y})\,\mathrm{d}\mathbf{y},$
$=\int[\nabla_{\mathbf{y}}\cdot\nabla_{\mathbf{y}}]\log p(\mathbf{y})p(\mathbf{y})\,\mathrm{d}\mathbf{y}+\frac{1}{2}\int\|\nabla_{\mathbf{y}}\log p(\mathbf{y})\|^{2}p(\mathbf{y})\,\mathrm{d}\mathbf{y}.$
Using the divergence theorem,
$\int[\nabla_{\mathbf{y}}\cdot\nabla_{\mathbf{y}}]\log p(\mathbf{y})p(\mathbf{y})\,\mathrm{d}\mathbf{y}$ $=\int\nabla_{\mathbf{y}}\cdot[\nabla_{\mathbf{y}}\log p(\mathbf{y})p(\mathbf{y})]\,\mathrm{d}\mathbf{y}-\int[\nabla_{\mathbf{y}}\log p(\mathbf{y})]\cdot[\nabla_{\mathbf{y}}p(\mathbf{y})]\,\mathrm{d}\mathbf{y}$
$=\oint\nabla_{\mathbf{y}}\log p(\mathbf{y})p(\mathbf{y})\,\mathrm{d}S(\mathbf{y})-\int[\nabla_{\mathbf{y}}\log p(\mathbf{y})]\cdot[\nabla_{\mathbf{y}}\log p(\mathbf{y})]p(\mathbf{y})\,\mathrm{d}\mathbf{y}$
$=0-\int\|\nabla_{\mathbf{y}}\log p(\mathbf{y})\|^{2}p(\mathbf{y})\,\mathrm{d}\mathbf{y}$
where the surface integral is zero because $\|\nabla_{\mathbf{y}}\log p(\mathbf{y})\|\to 0$ as $\mathbf{y}\to 0$. The conclusion follows.
The entropy $H$ is said to be twice differentiable at $\mathbb{P}\in\mathcal{P}$ if it is differentiable at $\mathbb{P}\in\mathcal{P}$ with gradient $\nabla_{\mathbb{P}}H$ and there exists a $\mathcal{P}\otimes\mathcal{P}$-integrable function $\nabla^{2}_{\mathbb{P}}H:\mathcal{Y}\times\mathcal{Y}\to\overline{\mathbb{R}}$ such that
$\lim_{\alpha\to 0^{+}}\frac{1}{\alpha^{2}}\Big{[}H((1-\alpha)\mathbb{P}+\alpha\mathbb{Q})-H(\mathbb{P})-\alpha\langle\nabla_{\mathbb{P}}H,\mathbb{Q}-\mathbb{P}\rangle\Big{]}\\ =\frac{1}{2}\langle\nabla^{2}_{\mathbb{P}}H(\mathbb{Q}-\mathbb{P}),\mathbb{Q}-\mathbb{P}\rangle\\ =\frac{1}{2}\iint\nabla^{2}_{\mathbb{P}}H(x,x^{\prime})\,\mathrm{d}(\mathbb{Q}-\mathbb{P})(x)\,\mathrm{d}(\mathbb{Q}-\mathbb{P})(x^{\prime}).$ (18)
Since the entropy $H$ is concave, the second derivative $\nabla^{2}_{\mathbb{P}}H$ is negative definite in the sense that the right-hand side of (18) is always less or equal to zero. This can be seen by using that
$H((1-\alpha)\mathbb{P}+\alpha\mathbb{Q})\leq H(\mathbb{P})+\langle\nabla_{\mathbb{P}}H,(1-\alpha)\mathbb{P}+\alpha\mathbb{Q}-\mathbb{P}\rangle$
due to the supergradient inequality (8).
Proof of Theorem 13.
If the entropy $H:\mathcal{P}\to\overline{\mathbb{R}}$ is twice-differentiable, then the function $h(t)=H(t\mathbb{Q}+(1-t)\mathbb{P})$ can be expressed as a 1st order Taylor series expansion around $t=0$ as $h(t)=h(0)+th^{\prime}(0)+r_{2}(t)$ where the remainder $r_{2}(t)$ can be expressed in the integral form as
$r_{2}(t)=\frac{1}{2}\left\langle\mathbb{Q}-\mathbb{P},\left[\int_{0}^{1}(t-u)\nabla_{u\mathbb{Q}+(1-u)\mathbb{P}}^{2}H\,\mathrm{d}u\right](\mathbb{Q}-\mathbb{P})\right\rangle.$ (19)
Evaluating the Taylor series and the remainder term (19) for $t=1$ leads to the desired result. ∎
Appendix B Proofs for Section 3.1
Proof of Theorem 15.
Sejdinovic et al. (2013, Proposition 20, Remark 21) imply that for all $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{h}$, $\iint h(x,y)\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{Q}(y)$ is finite. The entropy $H(\mathbb{P})$ is concave, since
$0\leq H(\lambda\mathbb{P}+(1-\lambda)\mathbb{Q})-\lambda H(\mathbb{P})-(1-\lambda)H(\mathbb{Q})=-\lambda(1-\lambda)\iint h(x,x^{\prime})\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(x)\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(x^{\prime})$ (20)
for every $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{h}$ and $\lambda\in[0,1]$, if $h$ is a conditionally negative definite kernel by taking $\mathbb{P},\mathbb{Q}$ to be discrete measures and using approximation arguments. Furthermore, strict concavity of $H$ (that is, equality in (20) if and only if $\lambda\in\{0,1\}$) coincides with $h$ being strongly conditionally negative definite. The conclusion follows by applying Theorem 12 to $H$, where the subgradient can be computed using (9). ∎
Proof of Theorem 16.
Suppose that the kernel score is proper. Then $h$ must be conditionally negative definite and Berg et al. (1984, Proposition 3.3.2) implies that there exists a Hilbert space $\mathcal{H}$, a subset $\{\psi_{y}\}_{y\in\mathcal{Y}}\subseteq\mathcal{H}$ and a function $f:\mathcal{Y}\to\mathbb{R}$, such that
$h(x,y)=\|\psi_{x}-\psi_{y}\|_{\mathcal{H}}^{2}+f(x)+f(y)$ (21)
for every $x,y\in\mathcal{Y}$. This implies the first equality in (13). The second equality is a straight forward computation, and the statement on strict propriety follows from the last expression of the divergence at (13). The converse follows since $h$ at (21) is conditinally negative definite. ∎
Remark 2.
Requiring $h$ to be non-negative does not lead to any loss of generality. Indeed, by (21), the entropy of any $h$ can be written as
$H(\mathbb{P})=\frac{1}{2}\iint h(x,y)\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{P}(y)=\frac{1}{2}\iint\|\psi_{x}-\psi_{y}\|_{\mathcal{H}}^{2}\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{P}(y)+\int f(x)\,\mathrm{d}\mathbb{P}(x)$
where the linear term $\int f(x)\,\mathrm{d}\mathbb{P}(x)$ is inconsequential and the kernels $h$ and the non-negative kernel $(x,y)\mapsto\|\psi_{x}-\psi_{y}\|_{\mathcal{H}}^{2}$ lead to the same kernel score.
For the proof of Theorem 19, we need the following definition. The supergradient map $\mathbb{P}\mapsto h_{\mathbb{P}}$ is weakly continuous if $\lim_{t\to 0+}\langle h_{\mathbb{P}+t\Delta\mathbb{P}},\mathbb{Q}\rangle=\langle h_{\mathbb{P}},\mathbb{Q}\rangle$ for $\mathbb{P},\mathbb{P}^{\prime},\mathbb{Q}\in\mathcal{P}$ where $\Delta\mathbb{P}=\mathbb{P}^{\prime}-\mathbb{P}$.
Proof of Theorem 19.
By Theorem 12, there exists a concave function $H:\mathcal{P}\to\mathbb{R}$ and supergradients $\{h_{\mathbb{P}}\}_{\mathbb{P}\in\mathcal{P}}$ such that
$S(\mathbb{P},y)=H(\mathbb{P})+\langle h_{\mathbb{P}},\delta_{y}-\mathbb{P}\rangle$
for every $y\in\mathcal{Y}$. For $\mathbb{P},\mathbb{Q}\in\mathcal{P}$, using $d(\mathbb{P},\mathbb{Q})=d(\mathbb{Q},\mathbb{P})$, we obtain
$H(\mathbb{Q})-H(\mathbb{P})=\frac{1}{2}\langle h_{\mathbb{P}}+h_{\mathbb{Q}},\mathbb{Q}-\mathbb{P}\rangle.$ (22)
By (22), the Gateaux derivative of $H$ at $\mathbb{Q}$ in the direction $\Delta\mathbb{Q}=\mathbb{Q}^{\prime}-\mathbb{Q}$, where $\mathbb{Q}^{\prime}\in\mathcal{P}$, exists and can be evaluated as
$\lim_{t\to 0}\frac{H(\mathbb{Q}+t\Delta\mathbb{Q})-H(\mathbb{Q})}{t}=\frac{1}{2}\lim_{t\to 0}~\langle h_{\mathbb{Q}}+h_{\mathbb{Q}+t\Delta\mathbb{Q}},\Delta\mathbb{Q}\rangle=\langle h_{\mathbb{Q}},\Delta\mathbb{Q}\rangle$
using the weak continuity of $\mathbb{P}\mapsto h_{\mathbb{P}}$. Evaluating the Gateaux derivative of both sides of (22) with respect at $\mathbb{Q}$ in the direction $\Delta\mathbb{Q}=\mathbb{Q}^{\prime}-\mathbb{Q}$ gives
$\langle h_{\mathbb{Q}},\Delta\mathbb{Q}\rangle$ $=\lim_{t\to 0}\frac{1}{2t}\Big{[}\langle h_{\mathbb{P}}+h_{\mathbb{Q}+t\Delta\mathbb{Q}},\mathbb{Q}+t\Delta\mathbb{Q}-\mathbb{P}\rangle-\langle h_{\mathbb{P}}+h_{\mathbb{Q}},\mathbb{Q}-\mathbb{P}\rangle\Big{]}$
$=\lim_{t\to 0}\frac{1}{2t}\Big{[}\langle h_{\mathbb{P}}+h_{\mathbb{Q}+t\Delta\mathbb{Q}},\mathbb{Q}-\mathbb{P}\rangle-\langle h_{\mathbb{P}}+h_{\mathbb{Q}},\mathbb{Q}-\mathbb{P}\rangle\Big{]}$
$+\,\lim_{t\to 0}\frac{t}{2t}\Big{[}\langle h_{\mathbb{P}}+h_{\mathbb{Q}+t\Delta\mathbb{Q}},\Delta\mathbb{Q}\rangle\Big{]}$
$=\lim_{t\to 0}\frac{1}{2t}\Big{[}\langle h_{\mathbb{Q}+t\Delta\mathbb{Q}}-h_{\mathbb{Q}},\mathbb{Q}-\mathbb{P}\rangle\Big{]}+\frac{1}{2}\Big{[}\langle h_{\mathbb{P}}+h_{\mathbb{Q}},\Delta\mathbb{Q}\rangle\Big{]}$
It follows that
$\lim_{t\to 0}\frac{1}{t}\Big{[}\langle h_{\mathbb{Q}+t\Delta\mathbb{Q}}-h_{\mathbb{Q}},\mathbb{Q}-\mathbb{P}\rangle\Big{]}=\langle h_{\mathbb{Q}}-h_{\mathbb{P}},\Delta\mathbb{Q}\rangle$ (23)
for every $\mathbb{P},\mathbb{Q},\mathbb{Q}^{\prime}\in\mathcal{P}$. By taking the difference of (23) with $\mathbb{P}=\mathbb{P}^{\prime}\in\mathcal{P}$ and with $\mathbb{P}\in\mathcal{P}$, we get
$\lim_{t\to 0}\frac{1}{t}\Big{[}\langle h_{\mathbb{Q}+t\Delta\mathbb{Q}}-h_{\mathbb{Q}},\Delta\mathbb{P}\rangle\Big{]}=\langle h_{\mathbb{P}^{\prime}}-h_{\mathbb{P}},\Delta\mathbb{Q}\rangle.$ (24)
for every $\mathbb{P},\mathbb{P}^{\prime},\mathbb{Q},\mathbb{Q}^{\prime}\in\mathcal{P}$ where $\Delta\mathbb{P}=\mathbb{P}^{\prime}-\mathbb{P}$. Using (24) on itself, we get
$\langle h_{\mathbb{P}^{\prime}}-h_{\mathbb{P}},\Delta\mathbb{Q}\rangle$ $=\lim_{t\to 0}\frac{1}{t}\langle h_{\mathbb{Q}+t\Delta\mathbb{Q}}-h_{\mathbb{Q}},\Delta\mathbb{P}\rangle$ By (24)
$=\lim_{t\to 0}\frac{1}{t}\left[\lim_{s\to 0}\frac{1}{s}\Big{[}\langle h_{\mathbb{P}+s\Delta\mathbb{P}}-h_{\mathbb{P}},\mathbb{Q}+t\Delta\mathbb{Q}-\mathbb{Q}\rangle\Big{]}\right]$ By (24)
$=\lim_{s\to 0}\frac{1}{s}\langle h_{\mathbb{P}+s\Delta\mathbb{P}}-h_{\mathbb{P}},\Delta\mathbb{Q}\rangle$
$=\langle h_{\mathbb{Q}^{\prime}}-h_{\mathbb{Q}},\Delta\mathbb{P}\rangle.$
Thus,
$\langle h_{\mathbb{P}^{\prime}}-h_{\mathbb{P}},\Delta\mathbb{Q}\rangle=\langle h_{\mathbb{Q}^{\prime}}-h_{\mathbb{Q}},\Delta\mathbb{P}\rangle$ (25)
for every $\mathbb{P},\mathbb{P}^{\prime},\mathbb{Q},\mathbb{Q}^{\prime}\in\mathcal{P}$.
Let $x_{0},y_{0}\in\mathcal{X}$. We expand the dual mappings into integrals twice as follows:
$\langle h_{\mathbb{P}^{\prime}}-h_{\mathbb{P}},\Delta\mathbb{Q}\rangle$ $=\int\Big{[}h_{\mathbb{P}^{\prime}}(y)-h_{\mathbb{P}}(y)\Big{]}\,\mathrm{d}(\Delta\mathbb{Q})(y)$
$=\int\Big{[}h_{\mathbb{P}^{\prime}}(y)-h_{\mathbb{P}}(y)-h_{\mathbb{P}^{\prime}}(y_{0})+h_{\mathbb{P}}(y_{0})\Big{]}\,\mathrm{d}(\Delta\mathbb{Q})(y)$
$=\int\langle h_{\mathbb{P}^{\prime}}-h_{\mathbb{P}},\delta_{y}-\delta_{y_{0}}\rangle\,\mathrm{d}(\Delta\mathbb{Q})(y)$
$=\int\langle h_{\delta_{y}}-h_{\delta_{y_{0}}},\Delta\mathbb{P}\rangle\,\mathrm{d}(\Delta\mathbb{Q})(y)$ By (25)
$=\iint\langle h_{\delta_{y}}-h_{\delta_{y_{0}}},\delta_{x}-\delta_{x_{0}}\rangle\,\mathrm{d}(\Delta\mathbb{P})(x)\,\mathrm{d}(\Delta\mathbb{Q})(y)$
$=\iint\langle h_{\delta_{y}},\delta_{x}\rangle\,\mathrm{d}(\Delta\mathbb{P})(x)\,\mathrm{d}(\Delta\mathbb{Q})(y).$
Thus,
$\langle h_{\mathbb{P}^{\prime}}-h_{\mathbb{P}},\Delta\mathbb{Q}\rangle=\int_{\mathcal{X}}\int_{\mathcal{X}}\langle h_{\delta_{y}},\delta_{x}\rangle\,\mathrm{d}(\Delta\mathbb{P})(x)\,\mathrm{d}(\Delta\mathbb{Q})(y)$
and by symmetry,
$\langle h_{\mathbb{Q}^{\prime}}-h_{\mathbb{Q}},\Delta\mathbb{P}\rangle=\int_{\mathcal{X}}\int_{\mathcal{X}}\langle h_{\delta_{y}},\delta_{x}\rangle\,\mathrm{d}(\Delta\mathbb{Q})(x)\,\mathrm{d}(\Delta\mathbb{P})(y).$
By (25) and the Fubini-Tonelli theorem, we derive the integral representation
$\langle h_{\mathbb{Q}^{\prime}}-h_{\mathbb{Q}},\Delta\mathbb{P}\rangle=\langle h_{\mathbb{P}^{\prime}}-h_{\mathbb{P}},\Delta\mathbb{Q}\rangle=\int_{\mathcal{X}}\int_{\mathcal{X}}K(x,y)\,\mathrm{d}(\Delta\mathbb{P})(x)\,\mathrm{d}(\Delta\mathbb{Q})(y)$ (26)
where $K(x,y)=\tfrac{1}{2}\left[\langle h_{\delta_{y}},\delta_{x}\rangle+\langle h_{\delta_{x}},\delta_{y}\rangle\right]=\tfrac{1}{2}[h_{\delta_{y}}(x)+h_{\delta_{x}}(y)]$ for $x,y\in\mathcal{Y}$. By definition, $K$ is symmetric. Due to the integral representation (26), we can write
$\int_{\mathcal{X}}\left[h_{\mathbb{P}^{\prime}}(y)-h_{\mathbb{P}}(y)-\int_{\mathcal{X}}K(x,y)\,\mathrm{d}(\Delta\mathbb{P})(x)\right]\,\mathrm{d}(\Delta\mathbb{Q})(y)=0$
for every $\mathbb{Q},\mathbb{Q}^{\prime}\in\mathcal{P}$. By taking $\mathbb{Q}=\delta_{y}$ and $\mathbb{Q}^{\prime}=\delta_{y^{\prime}}$ for $y,y^{\prime}\in\mathcal{X}$ with $y^{\prime}$ fixed, we can see that $h_{\mathbb{P}^{\prime}}(y)-h_{\mathbb{P}}(y)-\int_{\mathcal{X}}K(x,y)\,\mathrm{d}(\Delta\mathbb{P})(x)=c$ for every $y\in\mathcal{X}$ some constant $c\in\mathbb{R}$. By fixing $\mathbb{P}^{\prime}\in\mathcal{P}$, we have
[…]
for some constant function $C:\mathcal{X}\to\mathbb{R}$. It follows that
$\langle h_{\mathbb{P}},\mathbb{Q}\rangle-\int_{\mathcal{X}}\int_{\mathcal{X}}K(x,y)\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{Q}(y)=l(\mathbb{Q})$
for some linear function $l:\mathcal{P}\to\overline{\mathbb{R}}$. Here, a function $l:\mathcal{P}\to\overline{\mathbb{R}}$ is said to be linear if $l(\alpha\mathbb{P}+(1-\alpha)\mathbb{Q})=\alpha\cdot l(\mathbb{P})+(1-\alpha)\cdot l(\mathbb{Q})$ for every $\alpha\in[0,1]$ and $\mathbb{P},\mathbb{Q}\in\mathcal{P}$.
By substituting in (22), we get (using symmetry of $K$)
$2[H(\mathbb{Q})-H(\mathbb{P})]$ $=\langle h_{\mathbb{P}}+h_{\mathbb{Q}},\mathbb{Q}-\mathbb{P}\rangle$
$=\langle h_{\mathbb{P}},\mathbb{Q}\rangle-\langle h_{\mathbb{P}},\mathbb{P}\rangle+\langle h_{\mathbb{Q}},\mathbb{Q}\rangle-\langle h_{\mathbb{Q}},\mathbb{P}\rangle$
$=\int\int K(x,y)\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{Q}(y)-\int\int K(x,y)\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{P}(y)$
$\quad+\int\int K(x,y)\,\mathrm{d}\mathbb{Q}(x)\,\mathrm{d}\mathbb{Q}(y)-\int\int K(x,y)\,\mathrm{d}\mathbb{Q}(x)\,\mathrm{d}\mathbb{P}(y)$
$\quad+l(\mathbb{Q})-l(\mathbb{P})+l(\mathbb{Q})-l(\mathbb{P})$
$=\int\int K(x,y)\,\mathrm{d}\mathbb{Q}(x)\,\mathrm{d}\mathbb{Q}(y)+2l(\mathbb{Q})-\int\int K(x,y)\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{P}(y)-2l(\mathbb{P}).$
It follows that
$H(\mathbb{P})=\frac{1}{2}\int\int K(x,y)\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{P}(y)+l(\mathbb{P})+C$
for some $C\in\mathbb{R}$. The addition of a linear function to an entropy leads to the same score, so we might as well write
$H(\mathbb{P})=\frac{1}{2}\int\int K(x,y)\,\mathrm{d}\mathbb{P}(x)\,\mathrm{d}\mathbb{P}(y),$
and the claim is shown. ∎
Remark 3 (Why the quadratic score on $\mathbb{R}^{d}$ is not a kernel score.).
Assume to the contrary. Then there exists a kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ such that
$\int(p(\mathbf{y})-q(\mathbf{y}))^{2}\,\mathrm{d}\mathbf{y}=\int k(\mathbf{y},\mathbf{y}^{\prime})(p(\mathbf{y})-q(\mathbf{y}))(p(\mathbf{y}^{\prime})-q(\mathbf{y}^{\prime}))\,\mathrm{d}\mathbf{y}\,\mathrm{d}\mathbf{y}^{\prime}$
for all $p,q$. Since these are quadratic forms in $p-q$, we can use the parallelogram law to derive
$\int(p_{1}(\mathbf{y})-q_{1}(\mathbf{y}))(p_{2}(\mathbf{y})-q_{2}(\mathbf{y}))\,\mathrm{d}\mathbf{y}=\int k(\mathbf{y},\mathbf{y}^{\prime})(p_{1}(\mathbf{y})-q_{1}(\mathbf{y}))(p_{2}(\mathbf{y}^{\prime})-q_{2}(\mathbf{y}^{\prime}))\,\mathrm{d}\mathbf{y}\,\mathrm{d}\mathbf{y}^{\prime}$
for any densities $p_{1},p_{2},q_{1},q_{2}$ on $\mathbb{R}^{d}$. Let $A$ be a compact subset of $\mathbb{R}^{d}$ and $\{e_{j}\}_{j=1}^{\infty}$ be an orthonormal basis of the space of functions $f:A\to\mathbb{R}$ satisfying $\int_{A}f(\mathbf{y})\,\mathrm{d}\mathbf{y}=0$. By taking
$p_{1}(\mathbf{y})=\frac{\max(e_{i}(\mathbf{y}),0)}{2\int|e_{i}(\mathbf{y})|\,\mathrm{d}\mathbf{y}},\quad q_{1}(\mathbf{y})=\frac{\max(-e_{i}(\mathbf{y}),0)}{2\int|e_{i}(\mathbf{y})|\,\mathrm{d}\mathbf{y}},\quad p_{2}(\mathbf{y})=\frac{\max(e_{j}(\mathbf{y}),0)}{2\int|e_{j}(\mathbf{y})|\,\mathrm{d}\mathbf{y}},\quad q_{2}(\mathbf{y})=\frac{\max(-e_{j}(\mathbf{y}),0)}{2\int|e_{j}(\mathbf{y})|\,\mathrm{d}\mathbf{y}},$
we derive
$\delta_{ij}=\int_{A}e_{i}(\mathbf{y})e_{j}(\mathbf{y})\,\mathrm{d}\mathbf{y}=\int_{A\times A}k(\mathbf{y},\mathbf{y}^{\prime})e_{i}(\mathbf{y})e_{j}(\mathbf{y}^{\prime})\,\mathrm{d}\mathbf{y}\,\mathrm{d}\mathbf{y}^{\prime}$
for every $1\leq i,j<\infty$. Since $\{e_{i}(\mathbf{y})e_{j}(\mathbf{y}^{\prime})\}_{i,j=1}^{\infty}$ form a basis for square-integrable functions on $A\times A$, we have
$\int_{A\times A}|k(\mathbf{y},\mathbf{y}^{\prime})|^{2}\,\mathrm{d}\mathbf{y}\,\mathrm{d}\mathbf{y}^{\prime}=\lim_{N\to\infty}\sum_{i,j=1}^{N}\left[\int_{A\times A}k(\mathbf{y},\mathbf{y}^{\prime})e_{i}(\mathbf{y})e_{j}(\mathbf{y}^{\prime})\,\mathrm{d}\mathbf{y}\,\mathrm{d}\mathbf{y}^{\prime}\right]^{2}=\lim_{N\to\infty}N=\infty.$
By Theorem 16, $k$ can be chosen to be a negative-definite kernel of the form $k(\mathbf{y},\mathbf{y}^{\prime})=-\langle\psi_{\mathbf{y}},\psi_{\mathbf{y}^{\prime}}\rangle$ which means $|k(\mathbf{y},\mathbf{y}^{\prime})|^{2}\leq|k(\mathbf{y},\mathbf{y})||k(\mathbf{y}^{\prime},\mathbf{y}^{\prime})|$. It follows that the entropy of the uniform density $\mathbf{1}_{A}(\mathbf{y})/\int_{A}\,\mathrm{d}\mathbf{y}$ is given by
$\int_{A}\|\psi_{\mathbf{y}}\|^{2}\frac{\,\mathrm{d}\mathbf{y}}{\int_{A}\,\mathrm{d}\mathbf{y}}=\int_{A}|k(\mathbf{y},\mathbf{y})|\frac{\,\mathrm{d}\mathbf{y}}{\int_{A}\,\mathrm{d}\mathbf{y}}=\infty.$
This contradicts the fact that the entropy of the quadratic score is finite for the uniform density. The conclusion follows.
Proof of Theorem 20.
According to the Levy-Khinchin Theorem (Sasvári, 2013, Theorem 3.12.12) (in the special case for Hermitian real-valued $f$), a continuous function $h$ on $\mathbb{R}^{d}$ is conditionally negative definite if and only if there exist (a) a scalar $c\in\mathbb{R}$, (b) a vector $\mathbf{a}\in\mathbb{R}^{d}$, (c) a positive semidefinite matrix $\mathbf{B}\in\mathbb{R}^{d\times d}$, and (d) a $\sigma$-finite Borel measure $\mu$ satisfying $\mu(\{\mathbf{0}\})=0$, $\mu(A)=\mu(-A)$ for $A\subseteq\mathbb{R}^{d}$ and $\int_{\mathbb{R}^{d}}\min(1,\|\mathbf{x}\|^{2})\,\mathrm{d}\mu(\mathbf{x})<\infty$ such that
$h(\mathbf{x})=c+\langle\mathbf{x},\mathbf{B}\mathbf{x}\rangle-\int_{\mathbb{R}^{d}}\left[e^{i\langle\mathbf{x},\mathbf{u}\rangle}-1-\langle\mathbf{x},\mathbf{u}\rangle\mathbf{1}{\{\|\mathbf{u}\|\leq 1\}}\right]\,\mathrm{d}\mu(\mathbf{u}).$
Substituting in $S(\mathbb{P},\mathbf{x})=d(\mathbb{P},\delta_{\mathbf{x}})+\tfrac{1}{2}h(\mathbf{0})$ and $d(\mathbb{P},\mathbb{Q})$ where $d(\mathbb{P},\mathbb{Q})=-\frac{1}{2}\iint h(\mathbf{x}-\mathbf{y})\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(\mathbf{x})\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(\mathbf{y})$ gives the desired result. The special case for negative definite kernel follows from using Bochner’s Theorem (Sasvári, 2013, Theorem 1.7.4.) instead of Levy-Khinchine. ∎
Proof of Theorem 21.
The measure $\mu$ is supported on $\mathbb{R}^{d}\setminus\{\mathbf{0}\}$, which we can identify with $(0,\infty)\times\mathbb{S}^{d-1}$ via the transformations $r=\|\mathbf{u}\|\in(0,\infty)$ and $\sigma=\mathbf{u}/\|\mathbf{u}\|\in\mathbb{S}^{d-1}$. By Rohlin’s measure disintegration theorem (see Simmons, 2012, Theorem 6.4), we can write
$\,\mathrm{d}\mu(\mathbf{u})=\,\mathrm{d}\nu_{r}(\sigma)\,\mathrm{d}\rho(r)$
where $\{\nu_{r}\}$ are probability measures on $\mathbb{S}^{d-1}$ for $r\in(0,\infty)$ and $\rho$ is a $\sigma$-finite measure on (0,)(0$\infty$)( 0 , ∞ ). In order for $\int_{\mathbb{R}^{d}}\min(1,\|\mathbf{x}\|^{2})\,\mathrm{d}\mu(\mathbf{x})<\infty$ to hold, we must have $\int_{0}^{\infty}\min(1,r^{2})\,\mathrm{d}\rho(r)<\infty$.
  1. 1.
    Suppose that $S(\mathbb{P}_{c},c\mathbf{y})=c^{\alpha}S(\mathbb{P},\mathbf{y})$ for every $c>0$. By Theorem 20, we have
    0 0 $=(c^{2}-c^{\alpha})(\mathbf{y}-\mathbf{m}_{\mathbb{P}})^{\top}\mathbf{B}(\mathbf{y}-\mathbf{m}_{\mathbb{P}}),$ (27)
    0 0 $=c^{\alpha}\int_{\mathbb{R}^{d}}|e^{i\mathbf{u}^{\top}\mathbf{y}}-f_{\mathbb{P}}(\mathbf{u})|^{2}\,\mathrm{d}\mu(\mathbf{u})-\int_{\mathbb{R}^{d}}|e^{ic\mathbf{u}^{\top}\mathbf{y}}-f_{\mathbb{P}}(c\mathbf{u})|^{2}\,\mathrm{d}\mu(\mathbf{u}),$ (28)
    since the Lèvy-Khinchine decomposition is unique. By doing a change of variables in the second equation, we get
    $\int_{\mathbb{R}^{d}}|e^{i\mathbf{u}^{\top}\mathbf{y}}-f_{\mathbb{P}}(\mathbf{u})|^{2}\Big{[}c^{\alpha}\,\mathrm{d}\mu(\mathbf{u})-\,\mathrm{d}\mu(\mathbf{u}/c)\Big{]}=0.$
    For $\mathbb{P}=\delta_{\mathbf{0}}$, this gives (using the fact that $\mu$ is even)
    $\int_{\mathbb{R}^{d}}\Big{[}2-e^{i\mathbf{u}^{\top}\mathbf{y}}-e^{-i\mathbf{u}^{\top}\mathbf{y}}\Big{]}\Big{[}c^{\alpha}\,\mathrm{d}\mu(\mathbf{u})-\,\mathrm{d}\mu(\mathbf{u}/c)\Big{]}\\ =2\int_{\mathbb{R}^{d}}\Big{[}1-e^{i\mathbf{u}^{\top}\mathbf{y}}\Big{]}\Big{[}c^{\alpha}\,\mathrm{d}\mu(\mathbf{u})-\,\mathrm{d}\mu(\mathbf{u}/c)\Big{]}=0.$
    Thus,
    $\int_{\mathbb{R}^{d}}\Big{[}1-e^{i\mathbf{u}^{\top}\mathbf{y}}\Big{]}c^{\alpha}\,\mathrm{d}\mu(\mathbf{u})=\int_{\mathbb{R}^{d}}\Big{[}1-e^{i\mathbf{u}^{\top}\mathbf{y}}\Big{]}\,\mathrm{d}\mu(\mathbf{u}/c).$
    Again, using the evenness of $\mu$ we can write this as
    $\int_{\mathbb{R}^{d}}\Big{[}e^{i\mathbf{u}^{\top}\mathbf{y}}-1-\langle\mathbf{y},\mathbf{u}\rangle\mathbf{1}{\{\|\mathbf{u}\|\leq 1\}}\Big{]}c^{\alpha}\,\mathrm{d}\mu(\mathbf{u})=\int_{\mathbb{R}^{d}}\Big{[}e^{i\mathbf{u}^{\top}\mathbf{y}}-1-\langle\mathbf{y},\mathbf{u}\rangle\mathbf{1}{\{\|\mathbf{u}\|\leq 1\}}\Big{]}\,\mathrm{d}\mu(\mathbf{u}/c),$
    Since, this is a Lèvy-Khinchine decomposition, we have by uniqueness that the two Lèvy measures must be equal, which gives
    $\mu(cA)=\mu(A)/c^{\alpha}$ (29)
    for every Borel set $A\subseteq\mathbb{R}^{d}$. From (29) we get
    $\,\mathrm{d}\nu_{cr}(\sigma)\,\mathrm{d}\rho_{c}(r)=\frac{1}{c^{\alpha}}\,\mathrm{d}\nu_{r}(\sigma)\,\mathrm{d}\rho(r)$
    where $\rho_{c}(A)=\rho(cA)$ for Borel sets $A\subseteq(0,\infty)$. By integrating over $\sigma\in\mathbb{S}^{d-1}$, we get
    $\,\mathrm{d}\rho_{c}(r)=\frac{1}{c^{\alpha}}\,\mathrm{d}\rho(r)$ (30)
    which implies $\rho((0,cr))=\frac{1}{c^{\alpha}}\rho((0,r))$ for $c,r>0$. By fixing $r=1$ and replacing $c$ with $r$, we get $\rho((0,r))=\frac{1}{r^{\alpha}}\rho((0,1))$, which implies that $\rho$ is absolutely continuous with respect to the Lebesgue measure on (0,)(0$\infty$)( 0 , ∞ ) and
    $\,\mathrm{d}\rho(r)=\smash{\frac{\alpha\rho((0,1))}{r^{1+\alpha}}}\,\mathrm{d}r.$
    Furthermore, by (30), $\,\mathrm{d}\nu_{cr}(\sigma)=\,\mathrm{d}\nu_{r}(\sigma)$ which means that $\nu_{r}=\nu$ for $r>0$, where $\nu=\nu_{1}$. It follows that
    $\,\mathrm{d}\mu(\mathbf{u})=\,\mathrm{d}\nu(\sigma)\,\frac{\alpha\rho((0,1))}{r^{1+\alpha}}\,\,\mathrm{d}r$
    where we can write $\rho((0,1))=\mu(\mathbb{S}^{d-1}\times(0,1))$. In order for a nonzero measure $\rho$ to satisfy $\int_{0}^{\infty}\min(1,r^{2})\,\mathrm{d}\rho(r)<\infty$, we must have $\alpha\in(0,2)$.
    From (27) and (28), we thus have two possibilities: (a) $\alpha=2$ implying $\alpha\not\in(0,2)$ and thus, $\mu$ is zero everywhere and (b) $\alpha\neq 2$ implying $\mathbf{B}=\mathbf{0}$ and $\mu$ has the above form with $\alpha\in(0,2)$ (without ruling out the possibility that it is zero everywhere).
  2. 2.
    If $S(\mathbb{P}_{\mathbf{U}},\mathbf{U}\mathbf{y})=S(\mathbb{P},\mathbf{y})$, then
    0 0 $=(\mathbf{y}-\mathbf{m}_{\mathbb{P}})^{\top}[\mathbf{B}-\mathbf{U}^{\top}\mathbf{B}\mathbf{U}](\mathbf{y}-\mathbf{m}_{\mathbb{P}}),$
    0 0 $=\int_{\mathbb{R}^{d}}|e^{i\mathbf{u}^{\top}\mathbf{y}}-f_{\mathbb{P}}(\mathbf{u})|^{2}\,\mathrm{d}\mu(\mathbf{u})-\int_{\mathbb{R}^{d}}|e^{i(\mathbf{U}^{\top}\mathbf{u})^{\top}\mathbf{y}}-f_{\mathbb{P}}(\mathbf{U}^{\top}\mathbf{u})|^{2}\,\mathrm{d}\mu(\mathbf{u}).$
    for every rotation $\mathbf{U}\in\mathrm{SO}(d)$. Arguing as previously, it follows that $\mathbf{B}=c\mathbf{I}$ for some $c\geq 0$. Moreover, we get that
    $\,\mathrm{d}\nu_{r}(\mathbf{U}\sigma)\,\mathrm{d}\rho(r)=\,\mathrm{d}\nu_{r}(\sigma)\,\mathrm{d}\rho(r).$
    Integrating over $r$ reveals that $\nu_{r}$ is invariant under rotation, that is, $\,\mathrm{d}\nu_{r}(\mathbf{U}\sigma)=\,\mathrm{d}\nu_{r}(\sigma)$, implying $\,\mathrm{d}\nu_{r}(\sigma)=\Gamma(d/2)\,\mathrm{d}\sigma/(2\pi^{d/2})$.
The conclusion for energy scores and the CRPS follows from the form of their spectral measures once we combine the above results:
$\,\mathrm{d}\mu(\mathbf{u})=\alpha\mu(\mathbb{S}^{d-1}\times(0,1))\cdot\frac{\Gamma(d/2)}{2\pi^{d/2}}\cdot\frac{\,\mathrm{d}\sigma\cdot r^{d-1}\,\mathrm{d}r}{r^{d+\alpha}}\propto\frac{1}{\|\mathbf{u}\|^{d+\alpha}}\,\mathrm{d}\mathbf{u}.$